### v1

# Rotationally symmetric translating solitons of fully nonlinear extrinsic geometric flows: Classification and Applications

Abstract.
We study rotationally symmetric translators for fully nonlinear extrinsic geometric flows driven by a curvature function, and we establish the fine asymptotics of bowl-type evolutions and, when admissible, the construction and classification of catenoidal-type solutions, together with their asymptotic behavior. Under natural structural and convexity assumptions, we also prove rigidity and uniqueness results within appropriate classes of graphical translators of such curvature flows.
Key words and phrases:
Translating solitons; Fully nonlinear extrinsic flow; Maximum principles; Elliptic PDEs.
2010 Mathematics Subject Classification:
53A10, 53C21, 53C42, 53E10, 35J60, 58J70
1. Introduction
Translating solitons of extrinsic geometric flows whose normal velocity is given by a positive curvature function arise as canonical models for type II singularities, where curvature concentration behaves more irregularly than for spheres and cylinders during their deformation, see [chow2020extrinsic, lynch2023differential]. Moreover, they also serve as building blocks for the long-term behavior of these extrinsic curvature flows, see [rengaswami2023ancient].
These solitons evolve by translation along a unit direction under the deformation of the flow. In particular, rotationally symmetric models are of special interest due to the rigidity phenomena that they induce within certain families of curvature functions, see [cogo2023rotational] and references therein. For example, in the mean curvature flow, these models include the bowl soliton and catenoidal translators, which play a central role in the construction of $\Delta$-wing solutions, thus completing, together with grim-reaper cylinders, the classification of translating solitons that can be written as graphs on convex domains in space, and therefore ruling out some possible singular behavior for this flow in space, see [hoffman2019graphical].
The first goal of this article is to develop a unified framework for rotationally symmetric translators of $\gamma$-flows, where $\gamma(\lambda)$ is a curvature function which it is also the normal velocity of a family of evolving hypersurfaces. The class of curvature functions that we are interested in this article are positive or signed, $\alpha$-homogeneous and symmetric function that are increasing in each variable when they are evaluated in the principal curvatures of this family of hypersurfaces. Then, by using an implicit formulation of the translator equation in terms of the functions $x=g_{\pm}(y,\pm 1)$ associated with the level sets of $\tilde{\gamma}(x,y,\ldots,y)=\pm 1$, we establish the fine asymptotic expansion at infinity of bowl-type translators, see Theorems 4.1 and 4.8.
Our second main contribution is the construction and classification of rotationally symmetric catenoidal translators for suitable signed curvature functions (Section 5). For each $R>0$, we construct the catenoidal translator $W_{R}$ as a hypersurface of revolution generated by a curve with a single neck of radius $R$, and we show that, after removing a large ball, $W_{R}$ possesses two branches given by the graph of a function. Furthermore, we prove that its upper branch has the same bowl-type asymptotics as in the positive curvature case, while the lower branch exhibits either a bowl-type end or a different growth regime, depending on the behavior of $g_{-}(\cdot,-1)$ near the origin, see Theorem 5.1 and the pictures below.
Figure 1. Catenoidal translator whose curvature function is continuous at the origin. Figure prepared by Ignacio McManus.
Figure 2. Catenoidal translator whose curvature function satisfies $g_{-}(0,-1)=0$ and $\partial_{y}g_{-}(0,-1)<0$. Figure prepared by Ignacio McManus.
Using these rotationally symmetric models, we obtain geometric consequences for curvature functions that satisfy $\gamma(0,1,\ldots,1)>0$ and $\alpha>\frac{1}{3}$:
  • We show that any smooth, strictly convex entire $\gamma$-translator that is smoothly asymptotic to the bowl-type solution must coincide with it up to vertical translation, see Theorem 6.1.
  • We prove that there is no strictly convex translator contained inside a round vertical cylinder $\mathbb{S}^{n-1}\times\mathbb{R}$, see Corollary 6.8.
We organize the paper as follows. In Section 2, we introduce the geometric setting, notation, and structural assumptions on the curvature function. In Section 3, we develop the barrier method for the construction of the catenoidal examples. In Section 4, we derive the fine asymptotic behavior for bowl-type evolutions. In Section 5 we carry out the construction and classification of catenoidal translators. In Section 6 we prove the rigidity and uniqueness results within appropriate classes of graphical translators under the given structural hypotheses.
2. Preliminaries
In this section we lay the foundations for the article by defining the class of extrinsic curvature flows determined by a curvature function, either positive or signed, and by introducing the evolutions given by translation in a fixed unit direction, known as translating solitons. We also establish the framework for the rotationally symmetric models of these self-similar evolutions.
Firstly, given an immersed hypersurface $\Sigma_{0}=F_{0}(\Sigma)\subset{\mathbb{R}}^{n+1}$, which we assume smooth throughout the article unless stated otherwise, we recall some basic notions of extrinsic geometry in ${\mathbb{R}}^{n+1}$. The principal curvatures $\lambda(p)=(\lambda_{1}(p),\ldots,\lambda_{n}(p))$ of $p\in\Sigma_{0}$ are the eigenvalues of the shape operator $\mathcal{W}:T_{p}\Sigma_{0}\to T_{p}\Sigma_{0}$, defined at $p\in\Sigma_{0}$ by $\mathcal{W}(\vec{X})=-\nabla_{\vec{X}}\vec{\nu}$, where $\nu$ is a unit normal vector with a fixed orientation in ${\mathbb{R}}^{n+1}$ and $\nabla$ is the Euclidean connection.
2.1. The class of $\gamma$-flows
The class of extrinsic geometric flows that we consider consists of evolutions of a given initial hypersurface $\Sigma_{0}\subset\mathbb{R}^{n+1}$ described by an immersed family of hypersurfaces $\Sigma_{t}=F(\Sigma,t)$ whose normal velocity depends on the principal curvatures of $\Sigma_{t}$.
Definition 2.1.
Given a symmetric smooth function $\gamma:\Gamma\to(0,\infty)$ supported over a symmetric cone, we say that an immersed hypersurface $\Sigma_{0}=F_{0}(\Sigma)$ in $\mathbb{R}^{n+1}$ evolves under the $\gamma$-flow if there exists a one-parameter family of immersions $F:\Sigma\times[0,T)\to\mathbb{R}^{n+1}$ that solves the Cauchy problem
[…]
where $\lambda(x,t)=(\lambda_{1}(x,t),\ldots,\lambda_{n}(x,t))\in\Gamma$ is the principal curvature vector associated with $\vec{\nu}(x,t)$ on $\Sigma_{t}=F(\Sigma,t)$, and $\vec{\nu}(x,t)$ is a unit normal vector of $\Sigma_{t}$ in $\mathbb{R}^{n+1}$ with a fixed orientation.
Next we specify the class of curvature functions $\gamma$ that we will work with.
Definition 2.2.
Let $\Gamma\subset\mathbb{R}^{n}$ be an open symmetric cone [^1] Invariant under positive scaling, with the origin as a vertex. containing the positive cone $\Gamma_{+}=\{\lambda\in\mathbb{R}^{n}:\lambda_{i}>0\}$, and let $\alpha>0$. We say that a function $\gamma:\Gamma\to(0,\infty)$ is a positive $\alpha$-homogeneous curvature function if it satisfies:
  1. (1)
    $\gamma(\lambda)$ is smooth and symmetric, i.e., $\gamma(\lambda_{\sigma(1)},\ldots,\lambda_{\sigma(n)})=\gamma(\lambda_{1},\ldots,\lambda_{n})$ for every permutation $\sigma$ of $\{1,\ldots,n\}$.
  2. (2)
    $\gamma(\lambda)$ is strictly increasing in each coordinate, i.e., $\dfrac{\partial\gamma}{\partial\lambda_{i}}(\lambda)>0$ for every $i=1,\ldots,n$.
  3. (3)
    $\gamma(\lambda)$ is $\alpha$-homogeneous, i.e., $\gamma(c\lambda)=c^{\alpha}\gamma(\lambda)$ for every $c>0$.
This class of curvature functions includes, for $\alpha=1$, the following examples:
  • Mean curvature: $\gamma=H=\lambda_{1}+\ldots+\lambda_{n}$, supported in $\Gamma_{1}=\{\lambda\in\mathbb{R}^{n}:H>0\}$.
  • Gauss-Kronecker curvature: $\gamma=\sqrt[n]{K}=\sqrt[n]{\lambda_{1}\ldots\lambda_{n}}$, supported in $\Gamma_{+}$.
  • More generally, Hessian-type functions: $\gamma=\sqrt[k-l]{\dfrac{S_{k}}{S_{l}}}$ supported in the Gårding cone $\Gamma_{k}=\{\lambda\in\mathbb{R}^{n}:S_{i}(\lambda)>0\text{ for }i=1,\ldots,k\}$. Here $S_{k}(\lambda)=\sum\limits_{1\leq i_{1}<\ldots<i_{k}\leq n}\lambda_{i_{1}}\ldots\lambda_{i_{k}}$ denotes the elementary symmetric polynomial of order $k$ in $n$ variables, with the convention $S_{0}=1$ and $S_{k}=0$ for $k>n$.
  • The $k$-convexity functions: $\gamma=\left(\sum\limits_{1\leq i_{1}<\ldots<i_{k}\leq n}\dfrac{1}{\lambda_{i_{1}}+\ldots+\lambda_{i_{k}}}\right)^{-1}$ supported in $\Gamma_{k}^{\prime}=\{\lambda\in\mathbb{R}^{n}:\lambda_{i_{1}}+\ldots+\lambda_{i_{k}}>0\}$.
  • The $k$-norm functions: $\gamma(\lambda)=\sqrt[k]{\lambda_{1}^{k}+\ldots+\lambda_{n}^{k}}$ supported in the positive cone $\Gamma_{+}$.
We emphasize that the hypotheses defining the class of positive, 1-homogeneous curvature functions, together with second-order structural properties such as convexity, concavity, and inverse concavity, supported on the positive cone $\Gamma_{+}$, guarantee short-time existence and uniqueness [^2] Up to tangential diffeomorphisms of $\Sigma_{t}$. for strictly convex ($\lambda\in\Gamma_{+}$) compact initial data without boundary, see [chow2020extrinsic].
This subsection concludes with the definition of the curvature functions that we will use for the construction of the catenoidal solutions described in Section 5.
Definition 2.3.
We call a positive $\alpha$-homogeneous curvature function $\gamma:\overline{\Gamma}\to[0,\infty)$ signed if:
  1. (1)
    $\alpha=\frac{p}{q}\in\mathbb{Q}$ with $\gcd(p,2)=\gcd(q,2)=\gcd(p,q)=1$, and there exists a solution $(x_{0},y_{0},\ldots,y_{0})\in\{\gamma=0\}$ distinct from $(0,\ldots,0)$, $(0,\ldots,0,1)$, and $(0,1,\ldots,1)$ such that $\partial_{x}\gamma(x_{0},y_{0},\ldots,y_{0})>0$.
  2. (2)
    There is an open cone $\tilde{\Gamma}:=\{(x,y)\in\mathbb{R}^{2}:(x,y,\ldots,y)\in\Gamma\}$ and a continuous extension $\tilde{\gamma}:\tilde{\Gamma}_{+}\cup\tilde{\Gamma}_{-}\to\mathbb{R}$, where
    $\tilde{\Gamma}_{\pm}=\{(x,y)\in\tilde{\Gamma}:\gamma(x,y,\ldots,y)\gtrless 0\},$
    such that $\gamma$ is of class $C^{1}$ on each $\tilde{\Gamma}_{\pm}$.
  3. (3)
    $\gamma(cx,cy,\ldots,cy)=c^{\alpha}\gamma(x,y,\ldots,y)$ for all $c\in\mathbb{R}\setminus\{0\}$.
These hypotheses are motivated as follows. The homogeneity extends to rational exponents whose numerator and denominator are not divisible by $2$, which implies that the associated ODE inherits an even symmetry; that is, if $u$ solves the translating equation (3), then u(x)u(-x) is also a solution. Moreover, the condition $\partial_{x}\tilde{\gamma}(x_{0},y_{0})>0$ allows us to rewrite the related ODEs in an implicit form, namely $x=g(y,z)$ where $\tilde{\gamma}(x,y)=z$, which permits a classical ODE analysis.
2.2. Translators
We now introduce translating solitons. These are self-similar solutions to (1) of the form $\Sigma_{t}=\Sigma_{0}+t\vec{e}_{n+1}$. In particular, each time slice $\Sigma_{t}$ satisfies a local translator equation.
Definition 2.4.
We say that $\Sigma_{0}$ is a translating soliton in direction $\vec{e}_{n+1}\in\mathbb{R}^{n+1}$ of the $\gamma$-flow if, at every point of $\Sigma_{0}$, we have
(2) $\gamma(\lambda)=\langle\vec{\nu},\vec{e}_{n+1}\rangle.$
Before giving examples of translators, we introduce the notions of degeneracy and nondegeneracy for curvature functions.
Definition 2.5.
We say that a positive $\alpha$-homogeneous curvature function $\gamma:\Gamma\to(0,\infty)$ is $k$-nondegenerate if $(\underbrace{0,\ldots,0}_{k},1,\ldots,1)\in\Gamma$, and it is $k$-degenerate if $\gamma(\underbrace{0,\ldots,0}_{k},1,\ldots,1)=0$.
This definition is motivated by the behavior of the $\gamma$-flow on round generalized cylinder of the form $\mathbb{S}^{n-p}\times\mathbb{R}^{p}$ in ${\mathbb{R}}^{n+1}$. Under $k$-nondegeneracy, cylinders with $p<k$ contract in finite time to a $k$-hyperplane, whereas under $k$-degeneracy, cylinders with $p\geq k$ are stationary solutions of the $\gamma$-flow.
We now list examples of translating solitons for general curvature functions:
  • $\alpha$-grim-reaper cylinders. For (n1)(n-1)-nondegenerate, normalized $\alpha$-homogeneous positive curvature functions (i.e., $\gamma(0,\ldots,0,1)=1$), the family of translating cylinders, unique up to scaling and rotation, isometric to $C\times\mathbb{R}^{n-1}$, where $C$ is a one-dimensional $\alpha$-grim-reaper solving $\lambda^{\alpha}=\langle\nu,e_{2}\rangle$, see [bourni2022ancient, chou2001curve]. These cylinders are geodesically complete for $\alpha\leq 1$. Moreover, for $\alpha\in\left(\tfrac{1}{2},1\right]$, the generating curves are bounded on a finite interval, so the cylinders lie in slabs of finite width; whereas for $\alpha\in(0,\tfrac{1}{2})$, they are defined in all $\mathbb{R}$, yielding entire graphical solutions.
  • Bowl-type translators. Rotationally symmetric translating graphs of the form $(x,u(|x|))$ that satisfy a dichotomy: either the graph is entire, or it is defined only on a ball of radius $\sqrt[\alpha]{\gamma^{-1}(1,\ldots,1)}$. The dichotomy depends on the (non)degeneracy, the degree of homogeneity, and the asymptotics of the implicit function $x=g_{+}(y,1)$ coming from $\gamma(x,y,\ldots,y)=1$; see [rengaswami2024classification]. Moreover, for $1$-nondegenerate $\gamma$, we have $u(|x|)=|x|^{\alpha+1}+o\!\left(|x|^{\alpha+1}\right)$ as $|x|\to\infty$.
  • Catenoidal translators. This family consists of non-convex rotationally symmetric hypersurfaces parameterized by their positive distance to the translation axis, and they are singular along the plane $\{x_{n+1}=0\}$ in a large class of ambient spaces. These solutions are currently known only for the elementary symmetric polynomials, i.e., $\gamma=S_{k}$. As in the bowl-type case, we encounter a dichotomy: for $k=1,\ldots,n-1$, the catenoidal translators are rotational bi-graphs over the complement of a ball where each branch is asymptotic to a vertical translate of the bowl-type soliton [clutterbuck2007stability]. For $k=n$, the catenoidal family instead exhibits a conical singularity; we refer to [lima2025translators] for further details on the classification of these solutions .
We also remark that in the mean curvature case $\gamma=H$, there is a large variety of translating solitons, including the $\Delta$-wings family and several semi-graphical translators constructed and classified in [martin2024classification].
2.3. Rotationally symmetric translators
In this subsection we derive the equations satisfied by translating solitons that enjoy rotational symmetry, for the class of curvature functions considered in this article.
Definition 2.6.
Let $\mathtt{c}(s)=(r(s),u(s))\subset{\mathbb{R}}^{2}$ be a smooth regular curve, where $s$ is the arc-length parameter. We define the rotational hypersurface generated by $\mathtt{c}$ with respect to the $x_{n+1}$-axis by
$\Sigma(\mathtt{c})=\bigl\{(x^{\prime},x_{n+1})\in{\mathbb{R}}^{n+1}:(|x^{\prime}|,x_{n+1})=(r,u)\in\mathtt{c}\bigr\},$
where $|x^{\prime}|=\sqrt{x_{1}^{2}+\ldots+x_{n}^{2}}$.
Using cylindrical coordinates $\{x_{n+1},|x^{\prime}|,\vec{\omega}\}$ on ${\mathbb{R}}^{n+1}$, where $\vec{\omega}$ are spherical coordinates on $\mathbb{S}^{n-1}\subset{\mathbb{R}}^{n}$, we write a point $p\in\Sigma(\mathtt{c})$ as $p=r\vec{\omega}+u\vec{e}_{n+1}$. The unit normal vector at $p$ with fixed orientation can then be written as
$\nu(p)=u^{\prime}\vec{\omega}-r^{\prime}\vec{e}_{n+1}=\sin(\theta(s))\vec{\omega}-\cos(\theta(s))\vec{e}_{n+1},$
where $\theta^{\prime}(s)=\kappa=u^{\prime\prime}r^{\prime}-u^{\prime}r^{\prime\prime}$ is the curvature of $\mathtt{c}(s)$ in ${\mathbb{R}}^{2}$.
The principal curvatures at $p\in\Sigma(\mathtt{c})$ are given by
$\lambda_{1}(p)=\kappa,\qquad\lambda_{2}(p)=\ldots=\lambda_{n}(p)=\frac{\sin(\theta(s))}{r}.$
Consequently, the ODE system that characterizes a rotationally symmetric translator is
(3) $\begin{cases}\tilde{\gamma}\!\left(\kappa,\dfrac{\sin(\theta)}{r}\right)=\cos(\theta),\\[4.0pt] \theta^{\prime}=\kappa,\quad r^{\prime}=\cos(\theta),\quad u^{\prime}=\sin(\theta),\end{cases}$
where $\tilde{\gamma}:\tilde{\Gamma}\to(0,\infty)$ is defined by $\tilde{\gamma}=\gamma\circ\phi(x,y)$ with $\tilde{\Gamma}=\phi^{-1}(\Gamma)\subset{\mathbb{R}}^{2}$, and $\phi:{\mathbb{R}}^{2}\to{\mathbb{R}}^{n}$ is given by $\phi(x,y)=(x,y,\ldots,y)$.
2.4. Implicit translator equations.
We now reformulate equation (3) using the implicit function theorem applied to
(4) $\tilde{\gamma}(x,y)=z.$
Proposition 2.7.
Let $\gamma:\Gamma\to(0,\infty)$ be a positive $\alpha$-homogeneous curvature function. Then the level set
$\bigl\{(x,y,z)\in\tilde{\Gamma}\times(0,\infty):\tilde{\gamma}(x,y)=z\bigr\}$
can be described as the graph of $x=g(y,z)$, where $g_{+}(y,z)\in\mathcal{C}^{1}(U_{+},V_{+})$ is the implicit function of $\tilde{\gamma}(x,y)=z$ over open sets $U_{+}\subset{\mathbb{R}}^{2}$ and $V_{+}\subset{\mathbb{R}}$ given by:
  • If $\gamma(\lambda)$ is $1$-nondegenerate, then
    $U_{+}$ $=\Bigl\{(y,z)\in(0,\infty)^{2}:\frac{z}{\tilde{\gamma}(1,1)}<y^{\alpha}<\frac{z}{\tilde{\gamma}(0,1)}\Bigr\},\qquad V_{+}=(0,\infty).$
  • If $\gamma(\lambda)$ is $1$-degenerate, then
    $U_{+}$ $=\Bigl\{(y,z)\in(0,\infty)^{2}:\frac{z}{\tilde{\gamma}(1,1)}<y^{\alpha}\Bigr\},\qquad V_{+}=(0,\infty).$
Proof.
We consider $G:\tilde{\Gamma}\times(0,\infty)\to{\mathbb{R}}$ defined by $G(x,y,z)=\tilde{\gamma}(x,y)-z$. We note the following:
  1. (1)
    The point $(x,y,z)=(\tfrac{1}{2},1,\tilde{\gamma}(\tfrac{1}{2},1))$ lies on the level set $\tilde{\gamma}(x,y)=z$, and $x=\tfrac{1}{2}$ is the unique first coordinate for this choice because $\tilde{\gamma}$ is strictly increasing in each variable. [^3] Any point in the positive two-dimensional Gårding cone $\Gamma^{2}_{+}$ also works.
  2. (2)
    For each fixed $x$, the derivative with respect to $(y,z)$ is the nonzero vector
    $(\partial_{y}G,\partial_{z}G)=\left(\sum_{i=2}^{n}\frac{\partial\gamma}{\partial\lambda_{i}},-1\right),$
    which is nonsingular on $\tilde{\Gamma}\times{\mathbb{R}}$. Moreover, $\partial_{x}G=\frac{\partial\gamma}{\partial\lambda_{1}}>0$.
  3. (3)
    Let $(x_{n},y_{n},z_{n})\in V_{+}\times U_{+}$ satisfy $G(x_{n},y_{n},z_{n})=0$ and assume $(y_{n},z_{n})\to(y_{0},z_{0})\in U_{+}$. Then:
    • If $\gamma(\lambda)$ is $1$-nondegenerate, we have $\tilde{\gamma}(0,1)<\tilde{\gamma}\!\left(\frac{x_{n}}{y_{n}},1\right)<\tilde{\gamma}(1,1)$, which implies $0<x_{n}<y_{n}$. In particular, $(x_{n})$ is bounded, so it admits a convergent subsequence in $V_{+}$. If a subsequence converges to $x_{0}=0$, then $z_{0}=\tilde{\gamma}(0,y_{0})$ gives $\frac{z_{0}}{y_{0}^{\alpha}}=\tilde{\gamma}(0,1)$, contradicting $(y_{0},z_{0})\in U_{+}$.
    • If $\gamma(\lambda)$ is $1$-degenerate, the same estimate yields $0<x_{n}<y_{n}$, so $(x_{n})$ again has a convergent subsequence in $V_{+}$. If a subsequence converges to $x_{0}=0$, then $z_{0}=\tilde{\gamma}(0,y_{0})$ gives $\frac{z_{0}}{y_{0}^{\alpha}}=\tilde{\gamma}(0,1)=0$, again contradicting $(y_{0},z_{0})\in U_{+}$.
Since $V_{+}\times U_{+}$ is connected (indeed, convex), the hypotheses of the global implicit function theorem are satisfied (we refer the reader to [global_implicit_function_theorem, Thm. 3]) and we obtain a unique $g\in\mathcal{C}^{1}(U_{+},V_{+})$ such that $G(g(y,z),y,z)=0$ on $U_{+}$. ∎
We now establish an analogous result for signed curvature functions.
Proposition 2.8.
Let $\gamma\in\mathcal{C}^{1}(\Gamma)$ be a signed $1$-nondegenerate normalized (i.e., $\tilde{\gamma}(0,1)=1$) $\alpha$-homogeneous curvature function, or a $1$-degenerate $\alpha$-homogeneous curvature function. Then the level set
$\bigl\{(x,y)\in\tilde{\Gamma}_{-}\cap((0,\infty)\times(-\infty,0)):\tilde{\gamma}(x,y)=-1\bigr\}$
can be described as the graph of $x=g_{-}(y,-1)$, where $g\in\mathcal{C}^{1}(U_{-},V_{-})$ is the implicit function of $\tilde{\gamma}(x,y)=-1$ over open sets $U_{-}\subset{\mathbb{R}}$ and $V_{-}\subset{\mathbb{R}}$ given by
$U_{-}=\{y\in(-1,0):-1<\tilde{\gamma}(-1,1)y^{\alpha}\},\qquad V_{-}=(0,\infty).$
Proof.
As in the previous proof, we only need to verify the compactness property for the $x$-coordinate. Let $(x_{n},y_{n})\in V_{-}\times U_{-}$ be a sequence with $\tilde{\gamma}(x_{n},y_{n})=-1$ and $y_{n}\to y_{1}\in U_{-}$. Then $\tilde{\gamma}\!\left(\frac{x_{n}}{y_{n}},1\right)<\tilde{\gamma}(-1,1)$, so $x_{n}\in(0,|y_{n}|)$ and hence $(x_{n})$ is bounded. We may thus assume (passing to a subsequence) that $x_{n}\to x_{1}\in V_{-}$. If $x_{1}=0$, then
$1=\tilde{\gamma}(0,y_{1})=\begin{cases}0,&\text{if $\gamma$ is $1$-degenerate},\\ y_{1},&\text{if $\gamma$ is $1$-nondegenerate},\end{cases}$
which is impossible in either case. The global implicit function theorem used in the previous proof then applies. ∎
Proposition 2.9.
The function $g_{+}:U_{+}\to V_{+}$ is decreasing in $y$ and increasing in $z$, and satisfies
$c\,g_{+}(y,z)=g_{+}(cy,c^{\alpha}z)\quad\text{for every }c>0.$
Furthermore, $g_{-}(y,-1)$ is decreasing in $y$.
Proof.
Differentiating $\tilde{\gamma}(g_{+}(y,z),y)=z$ with respect to $y$ and $z$ yields
$\tilde{\gamma}_{x}(g_{+}(y,z),y)\,\partial_{y}g_{+}(y,z)+\tilde{\gamma}_{y}(g_{+}(y,z),y)$ $=0,$
$\tilde{\gamma}_{x}(g_{+}(y,z),y)\,\partial_{z}g_{+}(y,z)$ $=1.$
Thus
$\partial_{y}g_{+}(y,z)=-\frac{\tilde{\gamma}_{y}(g_{+}(y,z),y)}{\tilde{\gamma}_{x}(g_{+}(y,z),y)}<0,\qquad\partial_{z}g_{+}(y,z)=\frac{1}{\tilde{\gamma}_{x}(g_{+}(y,z),y)}>0.$
For the scaling property, let $c>0$ and observe that
$\tilde{\gamma}(g_{+}(cy,c^{\alpha}z),y)=c^{\alpha}z\quad\Longrightarrow\quad\tilde{\gamma}\!\left(\frac{g_{+}(cy,c^{\alpha}z)}{c},y\right)=z=\tilde{\gamma}(g_{+}(y,z),y).$
Since $\tilde{\gamma}$ is strictly increasing in its first variable, we conclude that $g_{+}(cy,c^{\alpha}z)=c\,g_{+}(y,z)$. The monotonicity of $g_{-}(\cdot,-1)$ follows by the same argument applied on $U_{-}$. ∎
As an example, we can compute $g_{\pm}$ explicitly for the $1$-homogeneous normalized Hessian quotient curvature functions
$\gamma(\lambda)=\sqrt[l-k]{\frac{S_{k}(0,1,\ldots,1)}{S_{l}(0,1,\ldots,1)}}\sqrt[k-l]{\frac{S_{k}(\lambda)}{S_{l}(\lambda)}},\qquad 0\leq l<k\leq n.$
In this case,
$\tilde{\gamma}(x,y)=\sqrt[k-l]{\frac{\binom{n-1}{l}}{\binom{n-1}{k}}}\,y\sqrt[k-l]{\frac{\binom{n-1}{k}y+\binom{n-1}{k-1}x}{\binom{n-1}{l}y+\binom{n-1}{l-1}x}},$
supported on a suitable cone $\tilde{\Gamma}_{k,\pm}$. Solving for $x=g_{\pm}(y,z)$ we obtain
$g_{+}(y,z)$ $=\frac{\binom{n-1}{k}y\left(z^{k-l}-y^{k-l}\right)}{\binom{n-1}{k-1}y^{k-l}-\dfrac{\binom{n-1}{k}\binom{n-1}{l-1}}{\binom{n-1}{l}}z^{k-l}},$
$g_{-}(y,-1)$ $=\frac{\binom{n-1}{k}y\left((-1)^{k-l}-y^{k-l}\right)}{\binom{n-1}{k-1}y^{k-l}-\dfrac{\binom{n-1}{k}\binom{n-1}{l-1}}{\binom{n-1}{l}}(-1)^{k-l}},$
defined on
$U_{+}$ $=\left\{(y,z)\in\mathbb{R}^{2}:\frac{z}{\sqrt[k-l]{\tfrac{n-l}{n-k}}}<y<z\right\},$
$U_{-}$ $=\left\{y\in(-1,0):-1<\sqrt[k-l]{\frac{(n-2k)(n-l)\binom{n-1}{l}}{(n-k)(n-2l)\binom{n-1}{k}}}\,y\right\}.$
This completes the preliminaries and the derivation of the implicit translator equations that we will use in the subsequent sections.
3. Barrier methods
In this section we develop the barrier method for rotationally symmetric translators whose generating curve $\mathtt{c}(s)$ is given as a vertical graph $(r,u(r))$ in $\mathbb{R}^{2}$.
Corollary 3.1.
Let $\alpha>0$ and let $\gamma$ be a positive or signed $\alpha$-homogeneous curvature function. Then rotationally symmetric vertical graphs satisfy
(5) $u^{\prime\prime}=(1+u^{\prime 2})^{\beta+1}g_{\pm}\left(\frac{u^{\prime}}{r(1+u^{\prime 2})^{\beta}},\pm 1\right),\qquad\beta=\frac{\alpha-1}{2\alpha}.$
Proof.
This is a straightforward computation from (3). ∎
The barrier method for (5) combines a comparison principle with suitable barriers that ensure the implicit functions $g_{\pm}(y,\pm 1)$ are well defined. It also provides first-order control for analyzing the behavior of solutions.
Definition 3.2.
We call a function $w(r)$ a supersolution (respectively, subsolution) to the initial value problem
(6) $\begin{cases}v^{\prime}=(1+v^{2})^{\beta+1}\,g_{\pm}\left(\dfrac{v}{r(1+v^{2})^{\beta}},\pm 1\right),\\[4.0pt] v(r_{0})=v_{0},\end{cases}$
if $w$ satisfies the differential inequality in (6) with $\geq$” (respectively, $\leq$”) in place of equality for all $r\geq r_{0}$, and the initial data obey $v_{0}\leq w_{0}=w(r_{0})$ (respectively, $v_{0}\geq w_{0}=w(r_{0})$). A function $v$ is a solution if it is both a subsolution and a supersolution.
We note that $g_{+}(y,1)$ extends continuously to the endpoints of $U_{+}$ by using (4). Indeed,
[…]
which implies
$g_{+}\left(\frac{1}{\sqrt[\alpha]{\tilde{\gamma}(1,1)}},1\right)=\frac{1}{\sqrt[\alpha]{\tilde{\gamma}(1,1)}},\qquad g_{+}\left(\frac{1}{\sqrt[\alpha]{\tilde{\gamma}(0,1)}},1\right)=0.$
Similarly, $g_{-}(y,-1)$ extends to the left endpoint of $U_{-}$ via (4), with value $g_{-}(-1,-1)=-1/\tilde{\gamma}(1,1)$, by homogeneity of $\tilde{\gamma}$. The additional quantity
(7) $\overline{m}_{0}=\frac{-1}{\sqrt[\alpha]{\tilde{\gamma}(-1,1)}}\in(-1,0)$
is meaningful only when $\tilde{\gamma}(-1,1)>0$, and in that case $g_{-}(\overline{m}_{0},-1)=-\overline{m}_{0}$.
Proposition 3.3.
Let $r_{0}\geq 0$ and suppose $v_{0}\leq w_{0}$ (respectively, $v_{0}\geq w_{0}$), with
$\frac{v_{0}}{r_{0}(1+v_{0}^{2})^{\beta}}\in\begin{cases}\left[\dfrac{1}{\sqrt[\alpha]{\tilde{\gamma}(1,1)}},\dfrac{1}{\sqrt[\alpha]{\tilde{\gamma}(0,1)}}\right],&\gamma\text{ positive, $1$-nondegenerate},\\[6.0pt] \left[\dfrac{1}{\sqrt[\alpha]{\tilde{\gamma}(1,1)}},\infty\right),&\gamma\text{ positive, $1$-degenerate},\\[6.0pt] [-1,0)\ \text{or}\ [\overline{m}_{0},0),&\gamma\text{ signed, $1$-nondegenerate},\end{cases}$
where $\overline{m}_{0}$ is as in (7). Assume $v,w\in\mathcal{C}^{1}([r_{0},\infty))$ are a solution and a supersolution (respectively, subsolution) to (6). Then $v\leq w$ (respectively, $v\geq w$) on $r_{0},\infty)$.
Proof.
We present the supersolution case; the subsolution case is analogous. Let $v,w\in\mathcal{C}^{1}([r_{0},\infty))$ be a solution and a supersolution of (6) with $v_{0}\leq w_{0}=w(r_{0})$. Define
$d(r)=\int_{v(r)}^{w(r)}\frac{1}{(1+t^{2})^{\beta+1}}\,dt.$
Since $w$ is a supersolution, we have $d(r_{0})\geq 0$. Suppose, for contradiction, that there exists $r_{1}>r_{0}$ with $v(r_{1})>w(r_{1})$, so $d(r_{1})<0$.
Set $r_{2}=\inf\{r\in(r_{0},r_{1}]:d(r)<0\}$. Then $d(r_{2})=0$ and $d(r)\leq 0$ on $r_{2},r_{1}]$, which implies $v(r)\geq w(r)$ there. By the mean value theorem, there exists $r_{3}\in(r_{2},r_{1})$ such that
$0>d^{\prime}(r_{3})$ $=\frac{w^{\prime}(r_{3})}{(1+w(r_{3})^{2})^{\beta+1}}-\frac{v^{\prime}(r_{3})}{(1+v(r_{3})^{2})^{\beta+1}}$
$\geq g_{\pm}\left(\frac{w(r_{3})}{r_{3}(1+w(r_{3})^{2})^{\beta}},\pm 1\right)-g_{\pm}\left(\frac{v(r_{3})}{r_{3}(1+v(r_{3})^{2})^{\beta}},\pm 1\right).$
On $r_{2},r_{1}]$ we have $v\geq w$. Since $\alpha>0$ implies $\beta<\tfrac{1}{2}$, the map $x\mapsto\frac{x}{(1+x^{2})^{\beta}}$ is strictly increasing. With $r_{3}>0$ we obtain
$\frac{w(r_{3})}{r_{3}(1+w(r_{3})^{2})^{\beta}}\leq\frac{v(r_{3})}{r_{3}(1+v(r_{3})^{2})^{\beta}}.$
Because $g_{\pm}$ is decreasing in its first argument, this yields
$g_{\pm}\left(\frac{w(r_{3})}{r_{3}(1+w(r_{3})^{2})^{\beta}},\pm 1\right)\geq g_{\pm}\left(\frac{v(r_{3})}{r_{3}(1+v(r_{3})^{2})^{\beta}},\pm 1\right),$
so $d^{\prime}(r_{3})\geq 0$, contradicting $d^{\prime}(r_{3})<0$. Hence no such $r_{1}$ exists and $v\leq w$ on $r_{0},\infty)$. ∎
For the signed setting, we now construct barriers that we will use later to derive existence, uniqueness, and growth estimates.
Proposition 3.4.
Let $r_{0}>0$ and $v_{0}$ be such that $\dfrac{v_{0}}{r_{0}(1+v_{0}^{2})^{\beta}}\in[-1,0)\cup[\overline{m}_{0},0)$, and assume $\beta<\tfrac{1}{2}$, where $\overline{m}_{0}$ is as in (7). For $\overline{m}\in[-1,0)\cup[\overline{m}_{0},0)$, let $w_{\overline{m}}(r)$ be the non-positive implicit function defined by
$\frac{w_{\overline{m}}(r)}{r(1+w_{\overline{m}}(r)^{2})^{\beta}}=\overline{m},\qquad w_{\overline{m}}(r_{0})=v_{0}.$
Then $w_{\overline{m}}$ is a subsolution of the IVP (6) whenever $\overline{m}_{0}\leq\overline{m}$.
Proof.
By the implicit function theorem, the negative branch $r\mapsto w_{\overline{m}}(r)$ is well defined and decreasing, with $w_{\overline{m}}(r_{0})=v_{0}$. A direct computation shows that
$w_{\overline{m}}^{\prime}-(1+w_{\overline{m}}^{2})^{\beta+1}\,g_{-}\left(\frac{w_{\overline{m}}}{r(1+w_{\overline{m}}^{2})^{\beta}},-1\right)$
has the same sign as
(8) $\frac{\overline{m}}{1+(1-2\beta)w_{\overline{m}}^{2}}-g_{-}(\overline{m},-1).$
Since $g_{-}(y,-1)$ is decreasing in $y$ and $g_{-}(\overline{m}_{0},-1)=-\overline{m}_{0}$, we have
$g_{-}(\overline{m},-1)\leq-\overline{m}_{0}\qquad\text{for }\overline{m}_{0}\leq\overline{m},$
which makes the expression in (8) nonpositive. Hence $w_{\overline{m}}$ is a subsolution. ∎
We now construct a supersolution for the signed case, which completes the well-posedness of (6) under the prescribed assumptions.
Proposition 3.5.
Let $r_{0},v_{0}$ be as in Proposition 3.4 and let $\beta\in(-1,\tfrac{1}{2})$ [^4] Equivalently, $\alpha>\tfrac{1}{3}$.. Assume furthermore that $g_{-}(0,-1)$ is either 0, negative, or satisfies $\lim\limits_{y\to 0^{-}}g_{-}(y,-1)=$-$\infty$, and set $b=\partial_{y}g_{-}(0,-1)<0$ when this derivative is finite, otherwise choose any $b<0$. For any $a>0$, define $\overline{w}(r)=-ar^{b}$ with initial condition $\overline{w}(r_{0})=v_{0}$. Then $\overline{w}$ is a supersolution. Moreover, if $g_{-}(0,-1)=0$ and $\partial_{y}g_{-}(0,-1)$ is finite, then
$v=\overline{w}+o(\overline{w})\quad\text{as }r\to\infty.$
Proof.
We first assume $g_{-}(0,-1)\leq 0$ and $b=\partial_{y}g_{-}(0,-1)<0$. With $a=-v_{0}r_{0}^{-b}>0$, set $t=\frac{ar^{b-1}}{(1+a^{2}r^{2b})^{\beta}}=\frac{-\overline{w}}{r(1+\overline{w}^{2})^{\beta}}\to 0^{+}$ as $r\to\infty$, since $b<0$.
Then, we compute
(9) $\overline{w}^{\prime}-(1+\overline{w}^{2})^{1+\beta}g_{-}\left(\frac{\overline{w}}{r(1+\overline{w}^{2})^{\beta}},-1\right)=\frac{-a}{r^{1-b}}\left(b-(1+\overline{w}^{2})\frac{g_{-}(-t,-1)}{-t}\right),$
to show that $\overline{w}$ is a supersolution. Indeed, we first notice that
$\frac{g_{-}(-t,-1)}{-t}=\frac{g_{-}(0,-1)}{-t}+\frac{g_{-}(-t,-1)-g_{-}(0,-1)}{-t},\mbox{ as }t\to 0^{+}.$
Then, if $g_{-}(0,-1)<0$ or $\lim\limits_{y\to 0^{-}}g_{-}(y,-1)=$-$\infty$ then $\frac{g_{-}(0,-1)}{-t}\to\infty$ as $t\to 0^{+}$, so the parentheses in (9) are eventually nonnegative for $r$ large enough. On the other hand, if $g_{-}(0,-1)=0$, then $\frac{g_{-}(-t,-1)-g_{-}(0,-1)}{-t}\to\partial_{y}g_{-}(0,-1)=b$, and the sign of the parentheses in (9) is governed by $b\,\overline{w}^{2}<0$. Thus the entire right-hand side of (9) is positive for all sufficiently large $r$, so $\overline{w}$ is a supersolution for large $r$.
For the asymptotic behavior, we assume $g_{-}(0,-1)=0$ and there exists $\varepsilon_{0}>0$ such that for every $R\gg 1$ there is $r\geq R$ with $\left|\frac{v(r)-\overline{w}(r)}{\overline{w}(r)}\right|\geq\varepsilon_{0}$. Then, since $\overline{w}$ is a supersolution, we observe that $v\leq(1+\varepsilon_{0})\overline{w}$ holds for $r$ large enough. Then, by using the monotonicity of $g_{-}$ and $b<0$, we estimate
$v^{\prime}(r)$ $\geq g_{-}\left(\frac{-(1+\varepsilon_{0})ar^{b-1}}{(1+(1+\varepsilon_{0})^{2}a^{2}r^{2b})^{\beta}},-1\right)$
$\geq-\partial_{y}g_{-}(0,-1)\,\frac{(1+\varepsilon_{0})ar^{b-1}}{(1+(1+\varepsilon_{0})^{2}a^{2}r^{2b})^{\beta}}$
$\geq\frac{|b|(1+\varepsilon_{0})}{(1+\delta)^{\beta}}ar^{b-1}$
for some $\delta>(1+\varepsilon_{0})^{1/\beta}-1$ and all $r\geq R(\delta)$ large enough. However, by integrating the above inequality we notice that
$\frac{1+\varepsilon_{0}}{(1+\delta)^{\beta}}\overline{w}(r)\leq v(r)\leq\overline{w}(r),$
which contradicts the choice of $\delta$. Hence, we obtain $v=\overline{w}+o(\overline{w})$ as $r\to\infty$. ∎
As an example of a curvature function that satisfies the hypotheses of this proposition, we consider $\gamma_{k}=Q_{k}=\frac{S_{k}}{S_{k-1}}$. In this case $g_{-}(y,-1)=\frac{\binom{n-1}{k}y\left(-1-y\right)}{\left(\binom{n-1}{k-1}y+\frac{\binom{n-1}{k}\binom{n-1}{k-2}}{\binom{n-1}{k-1}}\right)}$, and we notice that $\partial_{y}g_{-}(0,-1)=\frac{-(n-k+1)}{k-1}$. Since this number indicates the exponent in the growth order, we will observe in the following sections that for this curvature function the lower branch of the catenoidal translator $W_{R}$, has a growth order of $\frac{-a(R)(k-1)}{2(k+1)-n}r^{\frac{2(k+1)-n}{k-1}}$, which is negative when $k<\frac{n-2}{2}$, zero when $k=\frac{n-2}{2}$ is an integer, and positive otherwise.
Figure 3. $k=3$ the gorwth order is $r^{-1}$, $k=4$ the gorwth order is $\ln(r)$, and the growth order is $2r^{\frac{1}{2}}$
4. Fine Asymptotic of Bowl-type Solutions
In this section, we will derive by the fine asymptotic behaviors of bowl-type solutions at infinity based on the 1-(non)degeneracy assumptions.
4.1. The $1$-nondegenerate curvature function case.
In this subsection, we establish the asymptotic behavior at infinity of the initial value problem
(10) $\begin{cases}v^{\prime}=(1+v^{2})^{\beta+1}g_{+}\left(\dfrac{v}{r(1+v^{2})^{\beta}},1\right)\\ v(r_{0})=v_{0}.\end{cases}$
Theorem 4.1.
Let $\gamma:\Gamma\to(0,\infty)$ be a 1-nondegenerate, $\alpha$-homogeneous, normalized ($\tilde{\gamma}(0,1)=1$) positive curvature functions with $\alpha>\frac{1}{3}$, or equivalently, $\frac{\alpha-1}{2\alpha}=\beta\in\left(-1,\frac{1}{2}\right)$. Then, any solution to Eq. (10) with $r_{0},v_{0}$ as in Prop. 3.3 satisfies
(11) $v(r)=r^{\alpha}-\frac{a}{r^{\alpha}}+\frac{b}{r^{3\alpha}}+o\!\left(r^{-4\alpha}\right),\text{ as }r\to\infty,$
where the coefficients are given by
$a=$ $-\alpha\left(\dfrac{\alpha}{\partial_{y}g_{+}(1,1)}+\beta\right),$
$b=$ $\frac{2a\alpha^{2}-\partial_{y}g_{+}(1,1)\left[(1-2a)\beta\left(1+\beta(1-2a)\right)-2a^{2}\beta\right]}{2\alpha\partial_{y}g_{+}(1,1)(1-2\beta)}$
$+\frac{\partial_{y}g_{+}(1,1)\left(\frac{a}{\alpha}+\beta\right)(3\alpha-1)(1-2a)-\alpha\partial^{2}_{y}g_{+}(1,1)\left(\frac{a}{\alpha}+\beta\right)^{2}}{2\alpha\partial_{y}g_{+}(1,1)(1-2\beta)}.$
We employ a bootstrapping scheme similar to that used in [clutterbuck2007stability] to obtain the asymptotic expansion at infinity of the bowl-type solution for $\gamma=H$. To verify that the formulas coincide in the case $\gamma=H$, where $\alpha=1$ and $\beta=0$, we note that $g_{+}(1,1)=(n-1)(1-y)$, so $\partial g_{+}(1,1)=-(n-1)$, a=(n1)a=(n-1), and $\partial^{2}_{y}g_{+}(1,1)=0$. Therefore $b=\frac{(n-4)}{(n-1)^{2}}$, and the asymptotic behavior is
$v=r-\frac{1}{(n-1)r}+\frac{(n-4)}{(n-1)r^{3}}+o(r^{-4}),\mbox{ as }r\to\infty.$
This expression agrees with the one given in [clutterbuck2007stability] after rescaling the solution via $r\to\frac{r}{n-1}$, which corresponds to the normalization used in our computations.
Consequently, we divide our approach into several propositions, each deriving the next term in the asymptotic expansion of $v$. First, throughout this section we denote by $g(y)$ the implicit function $g_{+}(y,1):\left(\frac{1}{\sqrt[\alpha]{\tilde{\gamma}(1,1)}},1\right)\to(0,\infty)$ from Proposition 2.7, that is, we set $g(y)=g_{+}(y,1)$. Then $g$ satisfies
(12) $g(1)=0,\quad\text{and}\quad g^{\prime}(1)=-\frac{\tilde{\gamma}_{y}(0,1)}{\tilde{\gamma}_{x}(0,1)}=-\frac{\alpha}{\tilde{\gamma}x(0,1)}<0,$
since Euler’s theorem on homogeneous functions yields
$\alpha=\sum_{i=2}^{n}\frac{\partial\gamma}{\partial\lambda_{i}}(0,1,\ldots,1)=\tilde{\gamma}_{y}(0,1).$
The first step was established in [rengaswami2024classification] as part of our bootstrapping scheme.
Proposition 4.2.
The slope satisfies $v(r)=r^{\alpha}+o\!\left(r^{\alpha}\right)$ as $r\to\infty$.
We continue the bootstrapping by assuming the existence of a function $\varphi(r)=o(r^{\alpha})$ such that $v(r)=r^{\alpha}+\varphi(r)$ for all sufficiently large $r$. In fact, $\varphi(r)$ satisfies
$\varphi^{\prime}=\left(1+r^{2\alpha}\Big[1+\frac{\varphi}{r^{\alpha}}\Big]^{2}\right)^{\beta+1}g\left(\frac{r^{\alpha-1}\Big[1+\frac{\varphi}{r^{\alpha}}\Big]}{\left(1+r^{2\alpha}\Big[1+\frac{\varphi}{r^{\alpha}}\Big]^{2}\right)^{\beta}}\right)-\alpha r^{\alpha-1}.$
Proposition 4.3.
We have $\varphi(r)\to 0$ as $r\to\infty$.
Proof.
We assume, by contradiction, that there exists $\varepsilon_{0}>0$ such that for every $R>0$ there is $r=r(R)\geq R$ with $|\varphi(r)|\geq\varepsilon_{0}$. Since $\varphi(r)=o(r^{\alpha})$, for some $r=r(R)\geq R\gg 1$ we have either
(13) $-\varepsilon_{0}<\frac{\varphi(r)}{r^{\alpha}}\leq\frac{-\varepsilon_{0}}{r^{\alpha}},\qquad\text{or}\qquad\frac{\varepsilon_{0}}{r^{\alpha}}\leq\frac{\varphi(r)}{r^{\alpha}}<\varepsilon_{0}.$
Using that $g(y)$ is decreasing and that $y=\dfrac{r^{\alpha-1}x}{(1+r^{2\alpha}x^{2})^{\beta}}$ is increasing in $x$, together with $\beta>-1$, we obtain for the first case that
$\varphi^{\prime}$ $\geq\left(1+r^{2\alpha}\left[1-\varepsilon_{0}\right]^{2}\right)^{\beta+1}g\left(\frac{r^{\alpha-1}\left[1-\frac{\varepsilon_{0}}{r^{\alpha}}\right]}{\left(1+r^{2\alpha}\left[1-\frac{\varepsilon_{0}}{r^{\alpha}}\right]^{2}\right)^{\beta}}\right)-\alpha r^{\alpha-1},$
$\varphi^{\prime}$ $\leq\left(1+r^{2\alpha}\left[1+\varepsilon_{0}\right]^{2}\right)^{\beta+1}g\left(\frac{r^{\alpha-1}\left[1+\frac{\varepsilon_{0}}{r^{\alpha}}\right]}{\left(1+r^{2\alpha}\left[1+\frac{\varepsilon_{0}}{r^{\alpha}}\right]^{2}\right)^{\beta}}\right)-\alpha r^{\alpha-1},$
and similarly in the second case. Next, we compute the Laurent expansion at infinity of the auxiliary function
$f(r)=\frac{r^{\alpha-1}\big[1+\frac{\theta}{r^{\alpha}}\big]}{\left(1+r^{2\alpha}\big[1+\frac{\theta}{r^{\alpha}}\big]^{2}\right)^{\beta}}$
by setting $x=r^{-\alpha}$ and expanding at $x=0$,
$h(x)=f\big(x^{-1/\alpha}\big)=\frac{1+\theta x}{\big(x^{2}+[1+\theta x]^{2}\big)^{\beta}}=h(0)+h^{\prime}(0)x+o(x^{2}),$
where $\theta=\pm\varepsilon_{0}$, $h(0)=1$, and $h^{\prime}(0)=\theta(1-2\beta)=\dfrac{\theta}{\alpha}$. Hence the Laurent expansion of $f(r)$ at infinity is
$\frac{r^{\alpha-1}\big[1+\frac{\theta}{r^{\alpha}}\big]}{\left(1+r^{2\alpha}\big[1+\frac{\theta}{r^{\alpha}}\big]^{2}\right)^{\beta}}=1+\frac{\theta}{\alpha r^{\alpha}}+o(r^{-2\alpha}).$
Then the Taylor expansion of $g$ at $y=1$ gives
$g\left(\frac{r^{\alpha-1}\big[1+\frac{\theta}{r^{\alpha}}\big]}{\left(1+r^{2\alpha}\big[1+\frac{\theta}{r^{\alpha}}\big]^{2}\right)^{\beta}}\right)=\frac{g^{\prime}(1)\theta}{\alpha r^{\alpha}}+\text{l.o.t.}$
Therefore, by combining these expansions at infinity, we obtain either for the first case
$\varphi^{\prime}$ $\geq-(1-\varepsilon_{0})\varepsilon_{0}\frac{g^{\prime}(1)}{\alpha}r^{2\alpha-1}-\alpha r^{\alpha-1}+\mbox{ l.o.t.}>0,$
$\varphi^{\prime}$ $\leq(1+\varepsilon_{0})\varepsilon_{0}\frac{g^{\prime}(1)}{\alpha}r^{2\alpha-1}+\alpha r^{\alpha-1}+\mbox{ l.o.t.}<0,$
and for the second case. Then, we notice that, by continuity, either the first case forces the existence of a sequence $r_{k}\to\infty$ with $\varphi(r_{k})\geq Cr_{k}^{2\alpha}$, or the second case forces $\varphi(r_{k})\leq-Cr_{k}^{2\alpha}$, for some $C>0$. However, each alternative contradicts the corresponding estimate in (13). Therefore, no such $\varepsilon_{0}$ exists, which completes the proof. ∎
In what follows, to avoid making the proof overly tedious, whenever estimating the expansion at infinity of the next term in the asymptotic expansion of $v(r)$ yields a derivative bounded below or above by a multiple of $r^{2\alpha-1}$, we conclude the argument at that point and continue with the bootstrapping method.
Next, we introduce the function $\psi(r)$ by $\psi(r)=r^{\alpha}\varphi(r)\Leftrightarrow\varphi(r)=\frac{\psi(r)}{r^{\alpha}}$. This function $\psi(r)$ corresponds to the next term in the asymptotic expansion $v(r)=r^{\alpha}+\frac{\psi(r)}{r^{\alpha}}$. Moreover, we have $\psi(r)=o(r^{\alpha})$, and it satisfies
$\psi^{\prime}=r^{\alpha}\!\left(1+r^{2\alpha}\Big[1+\frac{\psi}{r^{2\alpha}}\Big]^{2}\right)^{\beta+1}\,g\!\left(\frac{r^{\alpha-1}\Big[1+\frac{\psi}{r^{2\alpha}}\Big]}{\left(1+r^{2\alpha}\Big[1+\frac{\psi}{r^{2\alpha}}\Big]^{2}\right)^{\beta}}\right)-\alpha r^{2\alpha-1}+\frac{\alpha\psi}{r}.$
Proposition 4.4.
Let $a=-\alpha\!\left(\dfrac{\alpha}{g^{\prime}(1)}+\beta\right)$. Then $\psi(r)\to-a$ as $r\to\infty$.
Proof.
With $a=-\alpha\!\left(\dfrac{\alpha}{g^{\prime}(1)}+\beta\right)$, we have
(14) $g^{\prime}(1)\Big[\tfrac{a}{\alpha}+\beta\Big]+\alpha=0.$
Proceeding as in the previous proposition, we assume there exists $\varepsilon_{0}>0$ such that for every $R>0$ there is $r=r(R)\geq R$ with $|\psi(r)+a|\geq\varepsilon_{0}$. Then, since $\psi=o(r^{\alpha})$, for some $r=r(R)\geq R\gg 1$ we have either
$-\frac{\varepsilon_{0}}{r^{\alpha}}<\frac{\psi}{r^{2\alpha}}\leq-\frac{\varepsilon_{0}+a}{r^{2\alpha}},\qquad\text{or}\qquad\frac{\varepsilon_{0}-a}{r^{2\alpha}}\leq\frac{\psi}{r^{2\alpha}}<\frac{\varepsilon_{0}}{r^{\alpha}}.$
Using the monotonicity of $g$ composed with $y=\dfrac{r^{\alpha-1}x}{(1+r^{2\alpha}x^{2})^{\beta}}$, we obtain either
$\psi^{\prime}\geq$ $\ r^{\alpha}\!\left(1+r^{2\alpha}\Big[1-\frac{\varepsilon_{0}}{r^{\alpha}}\Big]^{2}\right)^{\beta+1}\,g\!\left(\frac{r^{\alpha-1}\Big[1-\frac{\varepsilon_{0}+a}{r^{2\alpha}}\Big]}{\left(1+r^{2\alpha}\Big[1-\frac{\varepsilon_{0}+a}{r^{2\alpha}}\Big]^{2}\right)^{\beta}}\right)$
$-\alpha r^{2\alpha-1}-\alpha\varepsilon_{0}r^{\alpha-1},$
$\text{or }\quad\psi^{\prime}\leq$ $\ r^{\alpha}\!\left(1+r^{2\alpha}\Big[1+\frac{\varepsilon_{0}}{r^{\alpha}}\Big]^{2}\right)^{\beta+1}\,g\!\left(\frac{r^{\alpha-1}\Big[1+\frac{\varepsilon_{0}-a}{r^{2\alpha}}\Big]}{\left(1+r^{2\alpha}\Big[1+\frac{\varepsilon_{0}-a}{r^{2\alpha}}\Big]^{2}\right)^{\beta}}\right)$
$-\alpha r^{2\alpha-1}+\alpha\varepsilon_{0}r^{\alpha-1}.$
As before, we set
$f(r)=\frac{r^{\alpha-1}\left[1+\frac{\theta}{r^{2\alpha}}\right]}{\left(1+r^{2\alpha}\left[1+\frac{\theta}{r^{2\alpha}}\right]^{2}\right)^{\beta}},$
with $\theta\in\{-\varepsilon_{0}-a,\ \varepsilon_{0}-a\}$. With the substitution $r=x^{-1/(2\alpha)}$, we write
$h(x)=f\!\left(x^{-1/(2\alpha)}\right)=\frac{1+\theta x}{\big(x+[1+\theta x]^{2}\big)^{\beta}}=h(0)+h^{\prime}(0)\,x+o(x^{2}),$
where $h(0)=1$ and $h^{\prime}(0)=\theta(1-2\beta)-\beta=\dfrac{\theta}{\alpha}-\beta$. A Taylor expansion of $g$ at $y=1$ then yields
$g\!\left(\frac{r^{\alpha-1}\Big[1+\frac{\theta}{r^{2\alpha}}\Big]}{\left(1+r^{2\alpha}\Big[1+\frac{\theta}{r^{2\alpha}}\Big]^{2}\right)^{\beta}}\right)=-\,g^{\prime}(1)\left(\beta-\frac{\theta}{\alpha}\right)\frac{1}{r^{2\alpha}}+\text{l.o.t.}$
Since the leading contribution in
$r^{\alpha}\!\left(1+r^{2\alpha}\big[1\pm\frac{\varepsilon_{0}}{r^{\alpha}}\big]^{2}\right)^{\beta+1}$
is $r^{2\alpha}$, the resulting differential inequalities at infinity are
$\psi^{\prime}$ $\geq-\Big(g^{\prime}(1)\Big[\tfrac{a}{\alpha}+\beta\Big]+\alpha\Big)r^{2\alpha-1}-\frac{g^{\prime}(1)\,\varepsilon_{0}}{\alpha}\,r^{2\alpha-1}+\text{l.o.t.}=\,-\frac{g^{\prime}(1)\,\varepsilon_{0}}{\alpha}\,r^{2\alpha-1}+\text{l.o.t.},$
$\psi^{\prime}$ $\leq-\Big(g^{\prime}(1)\Big[\tfrac{a}{\alpha}+\beta\Big]+\alpha\Big)r^{2\alpha-1}+\frac{g^{\prime}(1)\,\varepsilon_{0}}{\alpha}\,r^{2\alpha-1}+\text{l.o.t.}=\ \frac{g^{\prime}(1)\,\varepsilon_{0}}{\alpha}\,r^{2\alpha-1}+\text{l.o.t.},$
where we use (14) in the equalities. In either case, integrating leads to a contradiction with $\psi=o(r^{\alpha})$ unless $\psi(r)\to-a$. Hence we conclude that $\psi(r)\to-a$ as $r\to\infty$. ∎
Now, to identify the next term in the asymptotic expansion of $v=r^{\alpha}+\dfrac{\psi}{r^{\alpha}}$, we consider the function $A(r)$ defined by $A=r^{\alpha}(\psi+a)\quad\Leftrightarrow\psi=\frac{A}{r^{\alpha}}-a$. As before, we have $A=o(r^{\alpha})$, and that it satisfies
$A^{\prime}=$ $\ r^{2\alpha}\left(1+r^{2\alpha}\Big[1+\frac{A}{r^{3\alpha}}-\frac{a}{r^{2\alpha}}\Big]^{2}\right)^{\beta+1}g\!\left(\frac{r^{\alpha-1}\Big[1+\frac{A}{r^{3\alpha}}-\frac{a}{r^{2\alpha}}\Big]}{\left(1+r^{2\alpha}\Big[1+\frac{A}{r^{3\alpha}}-\frac{a}{r^{2\alpha}}\Big]^{2}\right)^{\beta}}\right)$
$-\alpha r^{3\alpha-1}+2\alpha\,\frac{A}{r}-a\alpha r^{\alpha-1}.$
Proposition 4.5.
We have $A(r)\to 0$ as $r\to\infty$.
Proof.
As in the previous proofs, we assume there exists $\varepsilon_{0}>0$ such that for every $R\gg 1$ there is $r=r(R)\geq R$ with either
$-\frac{\varepsilon_{0}}{r^{2\alpha}}<\frac{A}{r^{3\alpha}}\leq-\frac{\varepsilon_{0}}{r^{3\alpha}},\qquad\text{or}\qquad\frac{\varepsilon_{0}}{r^{3\alpha}}\leq\frac{A}{r^{3\alpha}}<\frac{\varepsilon_{0}}{r^{2\alpha}}.$
By the monotonicity of each term in the equation for $A$, we obtain either
$A^{\prime}\geq{}$ $\ r^{2\alpha}\left(1+r^{2\alpha}\Big[1-\frac{\varepsilon_{0}+a}{r^{2\alpha}}\Big]^{2}\right)^{\beta+1}g\!\left(\frac{r^{\alpha-1}\Big[1-\frac{\varepsilon_{0}+ar^{\alpha}}{r^{3\alpha}}\Big]}{\left(1+r^{2\alpha}\Big[1-\frac{\varepsilon_{0}+ar^{\alpha}}{r^{3\alpha}}\Big]^{2}\right)^{\beta}}\right)$
$-\alpha r^{3\alpha-1}-(2\varepsilon_{0}+a)\alpha r^{\alpha-1},$
or
$A^{\prime}\leq{}$ $\ r^{2\alpha}\left(1+r^{2\alpha}\Big[1+\frac{\varepsilon_{0}-a}{r^{2\alpha}}\Big]^{2}\right)^{\beta+1}g\!\left(\frac{r^{\alpha-1}\Big[1+\frac{\varepsilon_{0}-ar^{\alpha}}{r^{3\alpha}}\Big]}{\left(1+r^{2\alpha}\Big[1+\frac{\varepsilon_{0}-ar^{\alpha}}{r^{3\alpha}}\Big]^{2}\right)^{\beta}}\right)$
$-\alpha r^{3\alpha-1}+(\,2\varepsilon_{0}-a\,)\alpha r^{\alpha-1}.$
Next, we consider
$f(r)=\frac{r^{\alpha-1}\left[1+\frac{\theta}{r^{3\alpha}}-\frac{a}{r^{2\alpha}}\right]}{\left(1+r^{2\alpha}\left[1+\frac{\theta}{r^{3\alpha}}-\frac{a}{r^{2\alpha}}\right]^{2}\right)^{\beta}},$
with $\theta\in\{-\varepsilon_{0},\ \varepsilon_{0}\}$. With the substitution $r=x^{-1/\alpha}$, we write
$h(x)=\frac{1+\theta x^{3}-ax^{2}}{\big(x^{2}+[1+\theta x^{3}-ax^{2}]^{2}\big)^{\beta}}=h(0)+h^{\prime}(0)\,x+\frac{h^{\prime\prime}(0)}{2}x^{2}+\frac{h^{\prime\prime\prime}(0)}{6}x^{3}+o(x^{4}),$
where $h(0)=1$, $h^{\prime}(0)=0$, $h^{\prime\prime}(0)=-2\!\left(\frac{a}{\alpha}+\beta\right)$, and $h^{\prime\prime\prime}(0)=6\theta(1-2\beta)=\dfrac{6\theta}{\alpha}$. A Taylor expansion of $g$ at $y=1$ yields
$g\!\left(\frac{r^{\alpha-1}\Big[1+\frac{\theta}{r^{3\alpha}}-\frac{a}{r^{2\alpha}}\Big]}{\left(1+r^{2\alpha}\Big[1+\frac{\theta}{r^{3\alpha}}-\frac{a}{r^{2\alpha}}\Big]^{2}\right)^{\beta}}\right)=-g^{\prime}(1)\left(\Big(\frac{a}{\alpha}+\beta\Big)\frac{1}{r^{2\alpha}}-\frac{\theta}{\alpha}\frac{1}{r^{3\alpha}}\right)+\text{l.o.t.}$
Finally, since the leading term in
$r^{2\alpha}\left(1+r^{2\alpha}\Big[1\pm\frac{\varepsilon_{0}}{r^{2\alpha}}-\frac{a}{r^{2\alpha}}\Big]^{2}\right)^{\beta+1}$
is $r^{5\alpha-1}$, and with our choice of $a$, we obtain at infinity
$A^{\prime}\geq-\,\frac{g^{\prime}(1)\,\varepsilon_{0}}{\alpha}\,r^{2\alpha-1}+\text{l.o.t.},\qquad\text{or}\qquad A^{\prime}\leq\ \frac{g^{\prime}(1)\,\varepsilon_{0}}{\alpha}\,r^{2\alpha-1}+\text{l.o.t.},$
which yields the desired contradiction unless $A(r)\to 0$. This completes the proof. ∎
For the next term, we introduce $B(r)=r^{\alpha}A(r)$. Then $B$ appears in $v(r)=r^{\alpha}-\dfrac{a}{r^{\alpha}}+\dfrac{B}{r^{3\alpha}}$ as $r\to\infty$. Moreover, we have $B(r)=o(r^{\alpha})$, and it satisfies
$B^{\prime}=$ $\ r^{3\alpha}\left(1+r^{2\alpha}\Big[1+\frac{B}{r^{4\alpha}}-\frac{a}{r^{2\alpha}}\Big]^{2}\right)^{\beta+1}g\!\left(\frac{r^{\alpha-1}\Big[1+\frac{B}{r^{4\alpha}}-\frac{a}{r^{2\alpha}}\Big]}{\left(1+r^{2\alpha}\Big[1+\frac{B}{r^{4\alpha}}-\frac{a}{r^{2\alpha}}\Big]^{2}\right)^{\beta}}\right)$
$-\alpha r^{4\alpha-1}-a\alpha r^{2\alpha-1}+3\alpha\,\frac{B}{r}.$
Proposition 4.6.
Let
$b=$ $\frac{2a\alpha^{2}-g^{\prime}(1)\left[(1-2a)\beta\left(1+\beta(1-2a)\right)-2a^{2}\beta\right]}{2\alpha g^{\prime}(1)(1-2\beta)}$
$+\frac{g^{\prime}(1)\left(\frac{a}{\alpha}+\beta\right)(3\alpha-1)(1-2a)-\alpha g^{\prime\prime}(1)\left(\frac{a}{\alpha}+\beta\right)^{2}}{2\alpha g^{\prime}(1)(1-2\beta)}.$
Then $B(r)\to b$ as $r\to\infty$.
Proof.
We assume there exists $\varepsilon_{0}>0$ such that for all $R\gg 1$ there is $r(R)\geq R$ with either
$-\frac{\varepsilon_{0}}{r^{3\alpha}}<\frac{B}{r^{4\alpha}}\leq-\frac{\varepsilon_{0}-b}{r^{4\alpha}},\qquad\text{or}\qquad\frac{\varepsilon_{0}+b}{r^{4\alpha}}\leq\frac{B}{r^{4\alpha}}<\frac{\varepsilon_{0}}{r^{3\alpha}}.$
By the monotonicity of the terms in the equation for $B$, we obtain
$B^{\prime}\geq{}$ $\ r^{3\alpha}\left(1+r^{2\alpha}\Big[1-\frac{\varepsilon_{0}+ar^{\alpha}}{r^{3\alpha}}\Big]^{2}\right)^{\beta+1}g\!\left(\frac{r^{\alpha-1}\Big[1-\frac{\varepsilon_{0}-b+ar^{2\alpha}}{r^{4\alpha}}\Big]}{\left(1+r^{2\alpha}\Big[1-\frac{\varepsilon_{0}-b+ar^{2\alpha}}{r^{4\alpha}}\Big]^{2}\right)^{\beta}}\right)$
$-\alpha r^{4\alpha-1}-a\alpha r^{2\alpha-1}-3\varepsilon_{0}r^{\alpha-1},$
or
$B^{\prime}\leq{}$ $\ r^{3\alpha}\left(1+r^{2\alpha}\Big[1+\frac{\varepsilon_{0}-ar^{\alpha}}{r^{3\alpha}}\Big]^{2}\right)^{\beta+1}g\!\left(\frac{r^{\alpha-1}\Big[1+\frac{\varepsilon_{0}+b-ar^{2\alpha}}{r^{4\alpha}}\Big]}{\left(1+r^{2\alpha}\Big[1+\frac{\varepsilon_{0}+b-ar^{2\alpha}}{r^{4\alpha}}\Big]^{2}\right)^{\beta}}\right)$
$-\alpha r^{4\alpha-1}-a\alpha r^{2\alpha-1}-3\varepsilon_{0}r^{\alpha-1}.$
We consider
$f(r)=\frac{r^{\alpha-1}\Big[1+\frac{\theta-ar^{2\alpha}}{r^{4\alpha}}\Big]}{\left(1+r^{2\alpha}\Big[1+\frac{\theta-ar^{2\alpha}}{r^{4\alpha}}\Big]^{2}\right)^{\beta}},\qquad\theta\in\{-\varepsilon_{0}+b,\ \varepsilon_{0}+b\}.$
Letting $r=x^{-1/(2\alpha)}$ gives
$h(x)=f(r)=\frac{1+\theta x^{2}-ax}{\big(x+[1+\theta x^{2}-ax]^{2}\big)^{\beta}}=h(0)+h^{\prime}(0)x+\frac{h^{\prime\prime}(0)}{2}x^{2}+o(x^{3}),$
where $h(0)=1$, $h^{\prime}(0)=-\big(\frac{a}{\alpha}+\beta\big)$, and
$h^{\prime\prime}(0)=2\theta(1-2\beta)+(1-2a)\beta\big(1+\beta(1-2a)\big)-2a^{2}\beta.$
A Taylor expansion of $g$ at $y=1$ yields
$g\!\left(f(r)\right)=g^{\prime}(1)\left(\frac{h^{\prime}(0)}{r^{2\alpha}}+\frac{h^{\prime\prime}(0)}{2r^{4\alpha}}\right)+\frac{g^{\prime\prime}(1)}{2}\left(\frac{h^{\prime}(0)}{r^{2\alpha}}+\frac{h^{\prime\prime}(0)}{2r^{4\alpha}}\right)^{2}+\text{l.o.t.}$
Finally, the term in
$r^{3\alpha}\left(1+r^{2\alpha}\Big[1\pm\frac{\varepsilon_{0}}{r^{3\alpha}}-\frac{a}{r^{2\alpha}}\Big]^{2}\right)^{\beta+1}$
that governs the bounds for $B^{\prime}$ is
$r^{6\alpha-1}+\frac{(3\alpha-1)}{2\alpha}(1-2a)\,r^{4\alpha-1},$
and, with our chosen constants $a$ and $b$, we obtain at infinity
$B^{\prime}\geq-\,\frac{g^{\prime}(1)\,\varepsilon_{0}}{\alpha}\,r^{2\alpha-1}+\text{l.o.t.},\qquad\text{or}\qquad B^{\prime}\leq\ \frac{g^{\prime}(1)\,\varepsilon_{0}}{\alpha}\,r^{2\alpha-1}+\text{l.o.t.},$
which implies that $B(r)\to b$ as $r\to\infty$. ∎
As a final step, we define $C=r^{\alpha}(B-b)$. Then $C$ appears in
$v(r)=r^{\alpha}-\dfrac{a}{r^{\alpha}}+\dfrac{b}{r^{3\alpha}}+\dfrac{C}{r^{4\alpha}},$
with $C(r)=o(r^{\alpha})$, and it satisfies
$C^{\prime}=$ $r^{4\alpha}\left(1+r^{2\alpha}\left[1+\frac{C}{r^{5\alpha}}+\frac{b}{r^{4\alpha}}-\frac{a}{r^{2\alpha}}\right]^{2}\right)^{\beta+1}$
$\times g\left(\frac{r^{\alpha-1}\left[1+\frac{C}{r^{5\alpha}}+\frac{b}{r^{4\alpha}}-\frac{a}{r^{2\alpha}}\right]}{\left(1+r^{2\alpha}\left[1+\frac{C}{r^{5\alpha}}+\frac{b}{r^{4\alpha}}-\frac{a}{r^{2\alpha}}\right]^{2}\right)^{\beta}}\right)$
$-\alpha r^{5\alpha-1}-a\alpha r^{3\alpha-1}+3\alpha b\,r^{\alpha-1}+4\alpha\,\frac{C}{r}.$
Proposition 4.7.
We have $C(r)\to 0$ as $r\to\infty$.
Proof.
We assume there exists $\varepsilon_{0}>0$ such that for all $R\gg 1$ there is $r=r(R)\geq R$ with either
$\frac{\varepsilon_{0}}{r^{4\alpha}}<\frac{C}{r^{5\alpha}}\leq-\frac{\varepsilon_{0}}{r^{5\alpha}},\qquad\text{or}\qquad\frac{\varepsilon_{0}}{r^{5\alpha}}\leq\frac{C}{r^{5\alpha}}<\frac{\varepsilon_{0}}{r^{4\alpha}}.$
By the monotonicity of the terms in the equation for $C$, we obtain
$C^{\prime}\geq{}$ $\ r^{4\alpha}\left(1+r^{2\alpha}\Big[1+\frac{b-\varepsilon_{0}-ar^{2\alpha}}{r^{4\alpha}}\Big]^{2}\right)^{\beta+1}g\!\left(\frac{r^{\alpha-1}\Big[1+\frac{br^{\alpha}-\varepsilon_{0}-ar^{3\alpha}}{r^{5\alpha}}\Big]}{\left(1+r^{2\alpha}\Big[1+\frac{br^{\alpha}-\varepsilon_{0}-ar^{3\alpha}}{r^{5\alpha}}\Big]^{2}\right)^{\beta}}\right)$
$-\alpha r^{5\alpha-1}+a\alpha r^{3\alpha-1}+3\alpha b\,r^{\alpha-1}-4\alpha\varepsilon_{0}r^{\alpha-1},$
or
$C^{\prime}\leq{}$ $\ r^{4\alpha}\left(1+r^{2\alpha}\Big[1+\frac{\varepsilon_{0}+b-ar^{2\alpha}}{r^{4\alpha}}\Big]^{2}\right)^{\beta+1}g\!\left(\frac{r^{\alpha-1}\Big[1+\frac{\varepsilon_{0}+br^{\alpha}-ar^{3\alpha}}{r^{5\alpha}}\Big]}{\left(1+r^{2\alpha}\Big[1+\frac{\varepsilon_{0}+br^{\alpha}-ar^{3\alpha}}{r^{5\alpha}}\Big]^{2}\right)^{\beta}}\right)$
$-\alpha r^{5\alpha-1}-a\alpha r^{3\alpha-1}-3\alpha b\,r^{\alpha-1}+4\alpha\varepsilon_{0}r^{\alpha-1}.$
We consider
$f(r)=\frac{r^{\alpha-1}\left[1+\frac{\theta+br^{\alpha}-ar^{3\alpha}}{r^{5\alpha}}\right]}{\left(1+r^{2\alpha}\left[1+\frac{\theta+br^{\alpha}-ar^{3\alpha}}{r^{5\alpha}}\right]^{2}\right)^{\beta}},\qquad\theta\in\{-\varepsilon_{0},\ \varepsilon_{0}\}.$
With $r=x^{-1/\alpha}$, we write
$h(x)$ $=\frac{1+\theta x^{5}+bx^{4}-ax^{2}}{\big(x^{2}+[1+\theta x^{5}+bx^{4}-ax^{2}]^{2}\big)^{\beta}}$
$=h(0)+h^{\prime}(0)x+\frac{h^{\prime\prime}(0)}{2}x^{2}+\frac{h^{\prime\prime\prime}(0)}{6}x^{3}+\frac{h^{(iv)}(0)}{24}x^{4}+\frac{h^{(v)}(0)}{120}x^{5}+o(x^{6}),$
where $h(0)=1$, $h^{\prime}(0)=0$, $h^{\prime\prime}(0)=-2\!\left(\frac{a}{\alpha}+\beta\right)$, $h^{\prime\prime\prime}(0)=0$, and
$h^{(iv)}(0)=24\!\left(b+a\beta(1-2a)-\beta(a^{2}+2b)+\frac{\beta(\beta+1)}{2}(1-2a)^{2}\right),\qquad h^{(v)}(0)=\frac{120\,\theta}{\alpha}.$
A Taylor expansion of $g$ at $y=1$ gives
$g\!\left(f(r)\right)$ $=g^{\prime}(1)\left(\frac{h^{\prime\prime}(0)}{2r^{2\alpha}}+\frac{h^{(iv)}(0)}{24r^{4\alpha}}+\frac{h^{(v)}(0)}{120}\frac{1}{r^{5\alpha}}\right)$
$\quad+\frac{g^{\prime\prime}(1)}{2}\left(\frac{h^{\prime}(0)}{r^{2\alpha}}+\frac{h^{\prime\prime}(0)}{2r^{4\alpha}}+\frac{h^{(v)}(0)}{120}\frac{1}{r^{5\alpha}}\right)^{2}+\text{l.o.t.}$
Finally, the term in
$r^{4\alpha}\left(1+r^{2\alpha}\Big[1\pm\frac{\varepsilon_{0}}{r^{3\alpha}}-\frac{a}{r^{2\alpha}}\Big]^{2}\right)^{\beta+1}$
that controls the bounds for $C^{\prime}$ is
$r^{7\alpha-1}+\frac{(3\alpha-1)}{2\alpha}(1-2a)\,r^{5\alpha-1},$
and, with our chosen constants $a$ and $b$, we obtain at infinity
$C^{\prime}\geq-\,\frac{g^{\prime}(1)\,\varepsilon_{0}}{\alpha}\,r^{2\alpha-1}+\text{l.o.t.},\qquad\text{or}\qquad C^{\prime}\leq\ \frac{g^{\prime}(1)\,\varepsilon_{0}}{\alpha}\,r^{2\alpha-1}+\text{l.o.t.},$
which implies $C(r)\to 0$ as $r\to\infty$. This completes the proof. ∎
By compiling the results of this section, we complete the proof of Theorem 4.1.
4.2. The $1$-degenerate curvature function case.
In this subsection, we establish the asymptotic behavior at infinity of entire bowl-type solutions to
(15) $\begin{cases}v^{\prime}=(1+v^{2})^{\beta+1}g_{+}\left(\dfrac{v}{r(1+v^{2})^{\beta}},1\right),\\ v(0)=0.\end{cases}$
for $1$-degenerate $\alpha$-homogeneous positive curvature functions. For instance, a condition ensuring the existence of entire bowl-type solutions for $1$-degenerate curvature functions is $\alpha\in\left(\frac{1}{3},\frac{1}{2}\right)$, or $g_{+}(y,1)\to 0$ with decaying growth $g_{+}(y,1)=O\!\left(y^{1-2\alpha}\right)$ as $y\to\infty$.
Theorem 4.8.
Let $\gamma:\overline{\Gamma}\to[0,\infty)$ be a $1$-degenerate $\alpha$-homogeneous positive curvature function with $k_{\gamma}>3\alpha-1$, and $\alpha\geq 1$. Then any solution to (15) behaves as $v=w+o(w^{-1})$ as $r\to\infty$ where $w(r)=A_{\gamma}r^{d_{\gamma}}$ and $d_{\gamma}=\frac{\alpha(k_{\gamma}+1)}{k_{\gamma}-2\alpha+1}$, $A_{\gamma}=\left(\frac{d_{\gamma}}{c_{\gamma}}\right)^{\frac{\alpha}{2\alpha-1-k_{\gamma}}}$, $k_{\gamma}=\min\{k\geq 1:\ c_{k}\neq 0\}$, where $c_{k}$ is the coefficient in the Laurent expansion at infinity of the implicit function $x=g_{+}(y,1)$, and $c_{\gamma}=c_{k_{\gamma}}>0$.
Remark 4.9.
Our assumption implies an estimate of the form
$g_{+}(y,1)=c_{\gamma}y^{-k_{\gamma}}+o\!\left(y^{-k_{\gamma}}\right)\qquad\text{as }y\to\infty$
for the implicit function from Proposition 2.8.
As in Section 4, we divide the proof of Theorem 4.8 into three propositions.
Proposition 4.10.
Let $\alpha\in\left[1,\infty\right)$, or equivalently $\beta\in\left(-\frac{1}{2},\frac{1}{2}\right)$, and assume that $k_{\gamma}\in(2\alpha-1,\infty)$. Then the function $w(r)$ from Theorem 4.8 is a subsolution to (15) if $c_{\gamma}>0$. Moreover, we have $v=w+o(w)$ as $r\to\infty$.
Proof.
A computation shows
$w^{\prime}-(1+w^{2})^{\beta+1}g_{+}\left(\frac{w}{r(1+w^{2})^{\beta}},1\right)$
$=$ $\,A_{\gamma}d_{\gamma}r^{d_{\gamma}-1}-(1+A_{\gamma}^{2}r^{2d_{\gamma}})^{\beta+1}g_{+}\left(\frac{A_{\gamma}r^{d_{\gamma}-1}}{(1+A_{\gamma}^{2}r^{2d_{\gamma}})^{\beta}},1\right)$
$=$ $\,A_{\gamma}d_{\gamma}r^{d_{\gamma}-1}-(1+A_{\gamma}^{2}r^{2d_{\gamma}})^{\beta+1}c_{\gamma}\left(\frac{(1+A_{\gamma}^{2}r^{2d_{\gamma}})^{\beta}}{A_{\gamma}r^{d_{\gamma}-1}}\right)^{k_{\gamma}}+\mbox{l.o.t.}$
$=$ $\,r^{d_{\gamma}-1}\underbrace{\left(A_{\gamma}d_{\gamma}-c_{\gamma}A_{\gamma}^{2(\beta+1)+(2\beta-1)k_{\gamma}}\right)}_{=0}$
$\ -c_{\gamma}\big(\beta(k_{\gamma}+1)+1\big)\,A_{\gamma}^{2(\beta(k_{\gamma}+1)+1)-k_{\gamma}-2}\,r^{\,2d_{\gamma}(\beta(k_{\gamma}+1)+1)-2d_{\gamma}-k_{\gamma}(d_{\gamma}-1)}+\text{l.o.t.}$
In the last identity, we use that
$d_{\gamma}-1=2(\beta+1)d_{\gamma}+d_{\gamma}k_{\gamma}(2\beta-1)+k_{\gamma}.$
Therefore, since $c_{\gamma}>0$, it follows that for $r$ large enough $w(r)$ is a subsolution to (15).
Next, we assume there exists $\varepsilon_{0}>0$ such that for every $R\gg 1$ there is $r=r(R)$ with $\left|\frac{v(r)-w(r)}{w(r)}\right|\geq\varepsilon_{0}$. We treat only the case $c_{\gamma}>0$, since $c_{\gamma}<0$ is analogous. Since $w(r)$ is a subsolution to (15), it follows that $v(r)\geq(1+\varepsilon_{0})w(r)$. Hence
$\dfrac{v^{\prime}}{(1+v^{2})^{\beta+1}}\leq g_{+}\left(\dfrac{w}{r(1+w^{2})^{\beta}},1\right)\leq c_{\gamma}A_{\gamma}^{\frac{-k_{\gamma}}{\alpha}}\,r^{k_{\gamma}\left(1-\frac{d_{\gamma}}{\alpha}\right)}.$
Now we set
$F(r)=\int_{v(R)}^{v(r)}\frac{1}{(1+t^{2})^{\beta+1}}\,dt,$
and we note that $F$ is increasing and satisfies $F(r)\leq v(r)$. The preceding inequality implies a bound of the form
$F(r)\leq F(R)-C_{\gamma}\left(r^{\frac{d_{\gamma}(1-2\alpha)}{\alpha}}-R^{\frac{d_{\gamma}(1-2\alpha)}{\alpha}}\right)$
for some constant $C_{\gamma}>0$. Since $\alpha>\frac{1}{2}$, we conclude that $v(r)$ is bounded by a constant depending on $F(R)$, $R$, and $C_{\gamma}$, which contradicts the fact that $w(r)$ is an increasing unbounded subsolution. ∎
To obtain the precise little-oh decay, we let $\varphi$ be such that $v(r)=w(r)+\varphi(r)$ and $\varphi=o(w(r))$ as $r\to\infty$. Then $\varphi$ satisfies
$\varphi^{\prime}=(1+(w+\varphi)^{2})^{\beta+1}g_{+}\left(\dfrac{w+\varphi}{r(1+[w+\varphi]^{2})^{\beta}},1\right)-A_{\gamma}d_{\gamma}r^{d_{\gamma}-1}.$
Proposition 4.11.
Under the assumptions of Proposition 4.10 together with $k_{\gamma}\geq 3\alpha-1$, we have $\varphi(r)\to 0$ as $r\to\infty$.
Proof.
Let $k_{\gamma}>3\alpha-1$ and assume there exists $\varepsilon_{0}>0$ such that for every $R\gg 1$ there is $r=r(R)\geq R$ with $|\varphi(r)|\geq\varepsilon_{0}$. Since $\varphi=o(w(r))$ as $r\to\infty$, for $R$ large enough we have either $\varepsilon_{0}\leq\varphi<Cw$ or $Cw<\varphi\leq-\varepsilon_{0}$, for some fixed $C>0$. In the first case, we obtain
$\varphi^{\prime}$ $\leq\left(1+(1+C)^{2}w^{2}\right)^{\beta+1}g_{+}\left(\frac{w+\varepsilon_{0}}{r\left(1+(1+C)^{2}w^{2}\right)^{\beta}},1\right)-A_{\gamma}d_{\gamma}r^{d_{\gamma}-1}$
$\leq A_{\gamma}d_{\gamma}r^{d_{\gamma}-1}\left[\left(1+C\right)^{\frac{3\alpha-1-k_{\gamma}}{\alpha}}-1+o(1)\right],$
while in the second case we analogously obtain
$\varphi^{\prime}$ $\geq\left(1+(1-C)^{2}w^{2}\right)^{\beta+1}g_{+}\left(\frac{w-\varepsilon_{0}}{r\left(1+(1-C)^{2}w^{2}\right)^{\beta}},1\right)-A_{\gamma}d_{\gamma}r^{d_{\gamma}-1}$
$\geq A_{\gamma}d_{\gamma}r^{d_{\gamma}-1}\left[\left(1-C\right)^{\frac{3\alpha-1-k_{\gamma}}{\alpha}}-1+o(1)\right].$
Since $k_{\gamma}>3\alpha-1$, it follows that $\varphi^{\prime}<0$ in the first case and $\varphi^{\prime}>0$ in the second case, up to an $o(1)$ error that does not change the sign for large $r$. This contradicts the monotonicity implied by the inequalities together with the lower bound in the first case and the upper bound in the second case. Therefore such an $\varepsilon_{0}$ cannot exist, and we conclude that $\varphi(r)\to 0$ as $r\to\infty$.
On the other hand, in the case $k_{\gamma}=3\alpha-1$, the asymptotic expansion obtained in the proof of Proposition 4.10 yields the precise sign needed to obtain an analogous contradiction. ∎
Proposition 4.12.
Under the hypotheses of Proposition 4.11, we have $\varphi=o(w^{-1})$.
Proof.
We assume there exist $\varepsilon_{0}>0$ and $C>0$ such that for all $R\gg 1$ there is $r=r(R)>R$ with $|\varphi(r)|\geq Cw$. In the case $Cw^{-1}\leq\varphi<\varepsilon_{0}$, we obtain
$\varphi^{\prime}$ $\geq\left(1+\left[w+\varepsilon_{0}\right]^{2}\right)^{\beta+1}g_{+}\left(\dfrac{w+\frac{C}{w}}{r\left(1+\left[w+\frac{C}{w}\right]^{2}\right)^{\beta}},1\right)-A_{\gamma}d_{\gamma}r^{d_{\gamma}-1}$
$\geq\frac{\varepsilon_{0}(3\alpha-1)}{\alpha r}+\mbox{l.o.t.}.$
This is impossible since it implies
$\varphi(r)\geq\varphi(R)+\frac{\varepsilon_{0}(3\alpha-1)}{\alpha}\ln\left(\frac{r}{R}\right),$
while $\varphi(r)\to 0$ as $r\to\infty$. The remaining case is analogous. ∎
The preceding proposition completes the proof of Theorem 4.8.
5. Catenoidal translators
In this section, we construct and classify the families of catenoidal translators $W_{R}$ for $1$-nondegenerate curvature functions.
Theorem 5.1.
Given a signed $\alpha$-homogeneous curvature function $\gamma:\overline{\Gamma}\to[0,\infty)$. Let $\alpha>\frac{1}{3}$ and $R>0$. Then we obtain a family of translating hypersurfaces of revolution $W_{R}$, called catenoidal translators, given by $\Sigma(\mathtt{c})$, where $\mathtt{c}(s)$ is a curve of class $\mathtt{c}\in\mathcal{C}^{2}\big((-\infty,0)\cup(0,\infty)\big)\cap\mathcal{C}^{1}\big((-\infty,\infty)\big)$ that satisfies Eq. (3) with initial data $\mathtt{c}(0)=(R,0)$ and $\mathtt{c}^{\prime}(0)=(0,1)$.
Furthermore, by removing the point $s=0$ we obtain two graphs,
$\mathtt{c}_{+}(r)$ $=(r,u_{+}(r)),\mbox{ with }r>0,\mbox{ and }$
$\mathtt{c}_{-}(\bar{r})$ $=(\bar{r},u_{-}(\bar{r})),\mbox{ with }\bar{r}=-r,$
defined on $(0,\infty)$, with the following asymptotic behavior:
(16) $u_{+}(r)=\frac{r^{\alpha+1}}{\alpha+1}+\frac{a}{(\alpha-1)\,r^{\alpha-1}}+o\!\left(r^{1-3\alpha}\right)+C_{+}(R)\quad\text{as }r\to\infty,$
for some $C_{+}(R)>0$. In addition, the asymptotic expansion of the lower branch depends on the following two alternatives:
  1. (1)
    $\tilde{\gamma}(0,0)=0$, with $\tilde{\gamma}$ continuous at $(0,0)$.
  2. (2)
    $g_{-}(0,-1)=0$ and $\partial_{y}g_{-}(0,-1)<0$, where $\tilde{\gamma}(g_{-}(y,-1),y)=-1$.
Then, in the first case,
(17) $u_{-}(r)=\frac{r^{\alpha+1}}{\alpha+1}+\frac{a}{(\alpha-1)\,r^{\alpha-1}}+o\!\left(r^{1-3\alpha}\right)+C_{-}(R)\quad\text{as }r\to\infty,$
for some $C_{-}(R)>0$. On the other hand, in the second case we have
$u_{-}(r)=f(r)+o\big(f(r)\big),\mbox{ where }f(r)=a\left(\frac{R^{b+1}}{b+1}-\frac{r^{b+1}}{b+1}\right),$
where $b=\partial_{y}g(0,-1)$ and $a=a(R)>0$.
First, by the initial conditions, we construct a small portion of the catenoidal translator as the “horizontal” graph of the distance function, i.e.,
$W_{R}\cap B(0,\varepsilon)=\bigcup_{u\in(-\varepsilon,\varepsilon)}r(u)\cdot\mathbb{S}^{n-1}\times\{u\},$
for some $\varepsilon=\varepsilon(R)>0$. Therefore, given a signed $\alpha$-homogeneous curvature function $\tilde{\gamma}:\tilde{\Gamma}\to\mathbb{R}$, we consider the following problem for this small portion in the variables $(r(u),u)$:
(18) $\begin{cases}\displaystyle\tilde{\gamma}\left(\dfrac{-r^{\prime\prime}}{(1+r^{\prime 2})^{1+\beta}},\dfrac{1}{r(1+r^{\prime 2})^{\beta}}\right)=r^{\prime},\\[6.0pt] r(0)=R,\quad r^{\prime}(0)=0.\end{cases}$
Proposition 5.2.
Given $R>0$, there exists $\varepsilon=\varepsilon(R)>0$ and a unique convex solution $r(u)\in\mathcal{C}^{2}((-\varepsilon,\varepsilon))$ to Eq. (18).
Proof.
First, the existence and uniqueness of a local solution to (18) follow from classical ODE theory. Indeed, since $\gamma(\lambda)$ is signed, it follows that $\partial_{x}\tilde{\gamma}(x_{0},y_{0})>0$ for some $(x_{0},y_{0})\notin\{(1,0),(0,0),(0,1)\}$. Hence, by the implicit function theorem there exist open neighborhoods $U\subset\mathbb{R}^{2}$ of $(y_{0},0)$ and $V\subset\mathbb{R}$ of $x_{0}$, with $V\times U\subset\overline{\tilde{\Gamma}}$, and a map $g:U\to V$ such that $\tilde{\gamma}(g(y,z),y)=z$.
Next, the function $r(u)$ satisfies
$r^{\prime\prime}=(1+r^{\prime 2})^{\beta+1}\,g\!\left(\frac{1}{r(1+r^{\prime 2})^{\beta}},\,r^{\prime}\right).$
Set $\rho=r^{\prime}$ and rewrite the problem as the first-order system
$\frac{d}{du}\begin{pmatrix}\rho\\ r\end{pmatrix}$ $=\begin{pmatrix}(1+\rho^{2})^{\beta+1}\,g\!\left(\dfrac{1}{r(1+\rho^{2})^{\beta}},\,\rho\right)\\[4.0pt] \rho\end{pmatrix}=F_{\gamma}(u,X).$
Since $g(y,z)$ has locally bounded partial derivatives on $U$, the vector field $F_{\gamma}(u,X)$ is locally Lipschitz. Therefore, by the Picard–Lindelöf theorem there exist $\varepsilon=\varepsilon(R)>0$ and a unique smooth solution $r(u)\in\mathcal{C}^{2}(-\varepsilon,\varepsilon)$ to the IVP (18).
Finally, possibly shrinking $\varepsilon$ if necessary, the solution is strictly convex, because $r^{\prime\prime}(0)=-g(R^{-1},0)>0$. Indeed,
$\tilde{\gamma}\left(R\,g(R^{-1},0),1\right)=0=\tilde{\gamma}\left(\frac{x_{0}}{y_{0}},1\right)$
is equivalent to $g(R^{-1},0)=\frac{x_{0}}{y_{0}\,R}<0$, since
$\frac{x_{0}}{y_{0}}=-\frac{\partial_{y}\tilde{\gamma}(x_{0},y_{0})}{\partial_{x}\tilde{\gamma}(x_{0},y_{0})}<0.$
This completes the proof. ∎
Next, we continue the construction of $W_{R}$ by superposing $W_{R}\cap B(0,\varepsilon)$ with the solutions of the IVP (6), namely $\mathtt{c}_{+}(s)=\mathtt{c}(s)$ for $s>0$ and $\mathtt{c}_{-}(s)=\mathtt{c}(s)$ for $s<0$. In addition, let $(s_{-},s_{+})$ be the maximal interval of existence of $\mathtt{c}(s)$, and define $\overline{\mathtt{c}}(s)=\mathtt{c}_{-}(s)$ on $(0,\overline{s})$, where $\overline{s}=-s_{-}$.
Proposition 5.3.
The angle function of $\mathtt{c}(s)$ satisfies $\theta(s)\in\left(0,\frac{\pi}{2}\right)$ and $\theta^{\prime}(s)>0$ for $s\in(0,s_{+})$.
Proof.
First, we assume that there exists a critical point $s_{0}\in(0,s_{+})$ of $\theta(s)$. Since $\theta^{\prime}(s)=\kappa(s)>0$ on $(0,s_{+})$, it follows that $\theta_{0}=\theta(s_{0})\in\left(0,\frac{\pi}{2}\right)$, and we set $r_{0}=r(s_{0})$. However, by computing the second derivative at the critical point via the implicit-function formulation,
$0\geq\left.\frac{d^{2}}{ds^{2}}\theta\right|_{s=s_{0}}=-\partial_{y}g_{+}\left(\frac{\sin(\theta_{0})}{r_{0}},\cos(\theta_{0})\right)\frac{\sin(2\theta_{0})}{2r_{0}^{2}}>0,$
we obtain a contradiction. Therefore, $\theta^{\prime}(s)>0$ for all $s\in(0,s_{+})$, and in particular $\theta(s)\in\bigl(0,\tfrac{\pi}{2}\bigr)$. ∎
As a consequence of the above proposition, we describe $\mathtt{c}_{+}(s)$ as the graph of a vertical function $u_{+}(r)$. Moreover, the slope $v(r)=u^{\prime}_{+}(r)$ satisfies Eq. (10) with initial condition $v(r_{0})=v_{0}$ as in Proposition 3.3. This implies that $s_{+}=\infty$, and the behavior at infinity of $u_{+}(r)$ corresponds to a primitive of Eq. (11) shifted by a positive constant $C_{+}(R)$, i.e., Eq. (16).
Now, we recall the setting for the lower branch of $W_{R}$:
$\overline{\mathtt{c}}(s)=\mathtt{c}_{-}(-s),\quad s\in(0,\overline{s}),\quad\overline{s}=-s_{-},\quad\overline{\mathtt{c}}(s)=(\overline{r}(s),\overline{u}(s)).$
We set $\overline{\theta}(s)=\theta(-s)+\pi$. Then, the following equations
$\overline{r}^{\prime}(s)=\cos(\overline{\theta}(s)),\quad\overline{u}^{\prime}(s)=\sin(\overline{\theta}(s))$
hold. Finally, we let $\overline{\kappa}(s)=\overline{\theta}^{\prime}(s)$; then $\overline{\theta}(s)\in\left(\frac{\pi}{2},\pi\right)$ for $s>0$ small enough. Thus $\overline{\mathtt{c}}(s)$ has the opposite orientation to $\mathtt{c}_{-}(s)$ [^5] This forces $\mathtt{c}(s)$ to be $\mathcal{C}^{2}$-singular at $s=0$. and it satisfies
(19) $\begin{cases}\displaystyle\tilde{\gamma}\left(\overline{\kappa},\dfrac{\sin(\overline{\theta})}{\overline{r}}\right)=\cos(\overline{\theta}),\qquad s\in(0,\overline{s}),\\[6.0pt] \overline{\mathtt{c}}(0)=(R,0),\quad\overline{\mathtt{c}}^{\prime}(0)=(0,-1).\end{cases}$
Proposition 5.4.
There exists $s_{0}\in(0,\overline{s})$ such that $\overline{\theta}(s_{0})=\frac{\pi}{2}$.
Proof.
We assume the opposite, i.e., $\overline{\theta}(s)\in\left(\frac{\pi}{2},\pi\right)$ for all $s\in(0,\overline{s})$. In particular,
$\overline{r}^{\prime}(s)=\cos(\overline{\theta}(s))<0,\quad\overline{u}^{\prime}(s)=\sin(\overline{\theta}(s))>0,$
and $\overline{\theta}(s)$ is strictly decreasing. Since $\overline{s}$ is the maximal time of existence, the limits $\overline{r}_{0}=\lim_{s\to\overline{s}}\overline{r}(s)\geq 0$, $\overline{u}_{0}=\lim_{s\to\overline{s}}\overline{u}(s)\in(0,\infty]$, and $\overline{\theta}_{0}=\lim_{s\to\overline{s}}\overline{\theta}(s)=\frac{\pi}{2}$ exist. Then, because $\overline{\kappa}(s)=\overline{\theta}^{\prime}(s)<0$, the curve $\overline{\mathtt{c}}(s)$ can be represented as the graph of a concave function $(\overline{r},\overline{u}(\overline{r}))$, with either $\overline{r}_{0}=0$ or $\lim_{\overline{r}\to\overline{r}_{0}\neq 0}\overline{u}(\overline{r})=\infty$.
We first consider the case $\lim\limits_{\overline{r}\to\overline{r}_{0}\neq 0}\overline{u}(\overline{r})=\infty$. Then there exists a sequence $s_{n}\in(0,\overline{s})$ with $s_{n}\to\overline{s}$ such that $\overline{r}(s_{n})\to\overline{r}_{0}>0$ and $\overline{\theta}(s_{n})\searrow\frac{\pi}{2}$. Evaluating (19) along this sequence gives
$\tilde{\gamma}\left(\overline{\kappa}(s_{n}),\frac{\sin(\overline{\theta}(s_{n}))}{\overline{r}(s_{n})}\right)=\cos(\overline{\theta}(s_{n})).$
Passing to the limit as $n\to\infty$ yields $\tilde{\gamma}\!\left(0,\frac{1}{\overline{r}_{0}}\right)=0$, in contradiction with the $1$-nondegeneracy assumption.
In the remaining case $\lim\limits_{\overline{r}\to 0}\overline{u}(\overline{r})=\infty$, we multiply (19) by $\overline{r}^{\alpha}(s_{n})$ and use the $\alpha$-homogeneity of $\tilde{\gamma}$; the same limit passage then leads again to a nonzero value on the left-hand side, whereas $\cos(\overline{\theta}(s_{n}))\to 0$, producing the same contradiction. Therefore, our assumption is false, and there exists $s_{0}\in(0,\overline{s})$ such that $\overline{\theta}(s_{0})=\frac{\pi}{2}$. ∎
Proposition 5.5.
Assume that $\tilde{\gamma}(x,y)$ has a continuous extension to the origin with $\tilde{\gamma}(0,0)=0$ and let $s_{0}$ be as in Proposition 5.4. Then, $\overline{\theta}$ has a global minimum at some $s_{1}\in(s_{0},\overline{s})$ and $\overline{\theta}^{\prime}>0$ on $(s_{1},\overline{s})$.
Proof.
We assume the opposite, i.e., $\overline{\theta}^{\prime}(s)<0$ on $(s_{0},\overline{s})$. Then, since now $\overline{r}^{\prime}=\cos(\overline{\theta})>0$ and $\overline{u}^{\prime}=\sin(\overline{\theta})>0$, the limits $\overline{r}_{\infty}=\lim_{s\to\overline{s}}\overline{r}(s)\in(R,\infty]$, $\overline{u}_{\infty}=\lim_{s\to\overline{s}}\overline{u}(s)=\infty$, and $\overline{\theta}_{\infty}=\lim_{s\to\overline{s}}\overline{\theta}(s)\in\left[0,\frac{\pi}{2}\right)$ exist.
Then $\overline{\mathtt{c}}(s)$ can be described as the graph of a concave function $(\overline{r},\overline{u}(\overline{r}))$. Therefore, we have $\overline{r}_{\infty}=\infty$, since otherwise the concave graph has a vertical asymptote to $\infty$.
As in the previous proof, there exists a sequence $s_{n}\to\overline{s}$ with $\overline{\theta}^{\prime}(s_{n})\to 0$ as $n\to\infty$. However, by evaluating at $s_{n}$ and taking limits in Eq. (19), we obtain $0=\tilde{\gamma}(0,0)=\cos(\overline{\theta}_{\infty})\in(0,1]$, which is impossible.
Consequently, there exists a first critical point $s_{1}>s_{0}$ of $\overline{\theta}$ with $\overline{\theta}(s_{1})\in\left(0,\frac{\pi}{2}\right)$. By computing the second derivative of $\overline{\theta}$ at $s_{1}$, we obtain
$\overline{\theta}^{\prime\prime}(s_{1})=-\partial_{y}g_{+}\left(\frac{\sin(\overline{\theta}(s_{1}))}{\overline{r}(s_{1})},\cos(\overline{\theta}(s_{1}))\right)\frac{\sin(2\overline{\theta}(s_{1}))}{2\overline{r}^{2}(s_{1})}>0.$
Hence $\overline{\theta}$ attains a strict minimum at $s_{1}$ and then increases.
As a summary, we describe $\overline{\mathtt{c}}(s)$ as the graph of a strictly convex function $(\overline{r},\overline{u}(\overline{r}))$ up to the next (possible) critical point $r_{2}=\overline{r}(s_{2})>\overline{r}(s_{1})$ of $\overline{\theta}$. However, since $\overline{\theta}(\overline{r})=\arctan\bigl(v(\overline{r})\bigr)$, with $v(\overline{r})=\overline{u}^{\prime}(\overline{r})$, it satisfies
$0$ $=\overline{\theta}^{\prime}(r_{2})=(1+v^{2})^{\beta}\,g_{+}\left(\frac{v}{r(1+v^{2})^{\beta}},1\right)$
$\geq g_{+}\!\left(\frac{v(r_{2})}{r_{2}(1+v^{2}(r_{2}))^{\beta}},1\right)-g_{+}(1,1)$
$=\partial_{y}g_{+}\!\left(\frac{v(r_{2})}{r_{2}(1+v^{2}(r_{2}))^{\beta}},1\right)\left(\frac{v(r_{2})}{r_{2}(1+v^{2}(r_{2}))^{\beta}}-1\right)$
$>0,$
since the barrier method provides a strict supersolution $w_{1}(\overline{r})$ with $m=1$ in the positive branch, see [rengaswami2024classification]. This is a contradiction. Consequently, we have $\overline{\theta}^{\prime}>0$ for $s\in(s_{1},\overline{s})$. ∎
As a corollary of the above proposition and of the barrier method, we obtain that $\overline{\mathtt{c}}(s)$ can be written as the graph of a strictly convex function $(\overline{r},\overline{u}(\overline{r}))$ with $\overline{r}\in(\overline{r}(s_{1}),\overline{r}(\overline{s}))$. Therefore, we have $\overline{r}(\overline{s})=\infty$ and $\overline{s}=\infty$, and there exists a positive constant $C_{-}(R)$ such that the asymptotic behavior of $\overline{u}(\overline{r})$ is given by Eq. (17) shifted by $C_{-}(R)$.
Proposition 5.6.
Under the assumption of Proposition 5.5, the curve $\mathtt{c}(s)$ with $s\in\mathbb{R}$ is embedded.
Proof.
Let $u_{\pm}(r)$ be the two branches of $\mathtt{c}(s)=(r(s),u(s))$ for $r\geq r_{1}=r(s_{1})$, where $s_{1}$ is from Proposition 5.5. We notice that $\mathtt{c}(s)$ is embedded for $s\in(-s_{1},s_{1})$ because on that interval it is concave. Next, following the proof of Proposition 3.3, we set
$d(r)=\int_{u_{-}^{\prime}(r)}^{u^{\prime}_{+}(r)}\frac{dt}{(1+t^{2})^{\beta+1}}.$
Since $u_{\pm}(r)$ are strictly convex graphs, the first intersection point $r_{2}$ is a first contact point, i.e.,
$u_{+}^{\prime}(r_{2})=u^{\prime}_{-}(r_{2})\quad\text{and}\quad d(r_{2})=0.$
However, this is impossible because
$0>d^{\prime}(r_{2})=g_{+}\left(\dfrac{u^{\prime}_{+}}{r(1+u^{\prime 2}_{+})^{\beta}},1\right)-g_{+}\left(\dfrac{u^{\prime}_{-}}{r(1+u^{\prime 2}_{-})^{\beta}},1\right)=0.$
This proves embeddedness. ∎
Finally, we consider the case where $\tilde{\gamma}(0,0)$ does not admit a continuous extension.
Proposition 5.7.
We assume that $\beta\geq 0$, or equivalently $\alpha\geq 1$, and let $s_{0}$ be as in Proposition 5.4, with $g_{-}(0,-1)=0$. Then $\lim\limits_{s\to\overline{s}}\overline{\theta}^{\prime}(s)=0$, and $\overline{s}=\infty$ whenever $\partial_{y}g_{-}(0,-1)<0$ exists.
Proof.
Recall that $\overline{\theta}(s_{0})=\frac{\pi}{2}$, and assume that $\overline{\theta}^{\prime}<0$ on $(s_{0},\overline{s})$. As in the previous argument, this assumption forces $\overline{\mathtt{c}}$ to be described by an increasing concave graph $(\overline{r},\overline{u}(\overline{r}))$ with $\overline{r}(s_{n})\to\infty$, $\overline{u}(s_{n})\to\infty$, and $\overline{\theta}(s_{n})\to\overline{\theta}_{0}\in\left[0,\frac{\pi}{2}\right)$ for some sequence $s_{n}\to\overline{s}$. In the present situation, we obtain
$\tilde{\gamma}(0,1)=1$ $=\tilde{\gamma}\left(\frac{\overline{\theta}^{\prime}(s_{n})}{\cos^{\frac{1}{\alpha}}(\overline{\theta}(s_{n}))},\frac{\sin(\overline{\theta}(s_{n}))}{\cos^{\frac{1}{\alpha}}(\overline{\theta}(s_{n}))\,\overline{r}(s_{n})}\right)$
$<\tilde{\gamma}\left(0,\frac{\sin(\overline{\theta}(s_{n}))}{\cos^{\frac{1}{\alpha}}(\overline{\theta}(s_{n}))\,\overline{r}(s_{n})}\right).$
However, this is impossible, since it implies
$\cos^{\frac{1}{\alpha}}(\overline{\theta}(s_{n}))\,\overline{r}(s_{n})<\sin(\overline{\theta}(s_{n})),$
and the right-hand side is bounded while the left-hand side blows up.
Finally, after rewriting the equation in terms of $v=u^{\prime}=-\overline{u}^{\prime}$, we obtain
$v^{\prime}=(1+v^{2})^{\beta+1}g_{-}\left(\frac{v}{r(1+v^{2})^{\beta}},-1\right).$
By choosing $a$ so that the initial condition of the slope function agrees with $v$, Proposition 3.5 yields $\overline{s}=\infty$ and
$v(r)=w(r)+o(\overline{w}(r)),$
where $\overline{w}(r)=-a\,r^{\partial_{y}g(0,-1)}$.
Finally, by combining the results of this section with the last proposition, we complete the proof of Theorem 5.1 by an integration procedure.
6. Uniqueness of rotationally symmetric graphs and Applications
In this section we prove the following uniqueness result.
Theorem 6.1.
Let $\alpha>\frac{1}{3}$ and let $\gamma:\overline{\Gamma}\to[0,\infty)$ be a $1$-nondegenerate, $\alpha$-homogeneous, normalized curvature function. Let $\Sigma$ be a smooth, strictly convex entire graphical $\gamma$-translator which is smoothly asymptotic to the bowl-type solution; that is, if $u_{\Sigma}:\mathbb{R}^{n}\to\mathbb{R}$ denotes the function whose graph defines $\Sigma$, then
$u_{\Sigma}(r)=\begin{cases}\displaystyle\frac{r^{\alpha+1}}{\alpha+1}+\frac{a}{(\alpha-1)\,r^{\alpha-1}}+o\bigl(r^{1-3\alpha}\bigr),&\alpha\neq 1,\\[6.0pt] \displaystyle\frac{r^{2}}{2}-a\ln r+o\bigl(r^{-2}\bigr),&\alpha=1,\end{cases}\qquad\text{as }r=|x|\to\infty,$
where $|x|=\sqrt{x_{1}^{2}+\cdots+x_{n}^{2}}$. Then $\Sigma$ coincides, up to a vertical translation, with the bowl-type solution. Equivalently, $u_{\Sigma}$ is rotationally symmetric and therefore unique up to vertical translations.
The proof relies on Alexandrov?s moving-plane method. In this argument, the strict convexity assumption $\Sigma\subset\mathbb{R}^{n+1}$ is crucial, since it allows us to reflect $\Sigma$ across planes without requiring that the curvature function $\gamma(\lambda)$ be continuously defined at $(0,\ldots,0)$. Moreover, for the method to work, we need a comparison principle; see Theorem 6.2 below from [Yo, Thm 1.2].
Theorem 6.2.
Let $\gamma:\overline{\Gamma}\to[0,\infty)$ be a positive curvature function, and let $\Sigma_{1},\Sigma_{2}\subset\mathbb{R}^{n+1}$ be two complete, embedded, connected $\gamma$-translators such that $\Sigma_{1}$ is strictly convex, and $\Sigma_{2}$ is convex. Then:
  1. (1)
    If there exists an interior point $p\in\Sigma_{1}\cap\Sigma_{2}$ such that the tangent spaces coincide at $p$, and if $\Sigma_{1}$ lies on one side of $\Sigma_{2}$ near $p$, then $\Sigma_{1}=\Sigma_{2}$.
  2. (2)
    If the boundaries $\partial\Sigma_{i}$ lie in the same hyperplane $\Pi$ and the intersections $\Sigma_{i}\cap\Pi$ are transversal, and if $\Sigma_{1}$ lies on one side of $\Sigma_{2}$ and there exists $p\in\partial\Sigma_{1}\cap\partial\Sigma_{2}$ such that the tangent spaces of $\Sigma_{i}$ and their boundaries coincide at $p$, then $\Sigma_{1}=\Sigma_{2}$.
Consequently, we use a one-parameter family of hyperplanes to reflect $\Sigma$ across them, and the comparison principle characterizes the possible first-contact points (in the interior, on the boundary, or at infinity). At any such contact, the rigidity coming from the asymptotic behavior of $\Sigma$ at infinity forces $\Sigma$ to be symmetric with respect to the corresponding plane through the origin. Finally, since equation (2) is invariant under rotations around the $x_{n+1}$-axis, the rotational symmetry of $\Sigma$ follows.
We now introduce the family of moving planes and the associated notation.
  • For $t\in\mathbb{R}$, we define the moving planes
    $\Pi_{t}=\{x\in\mathbb{R}^{n+1}:\mathbf{p}(x)=t\},\qquad\mathbf{p}(x)=x_{1}.$
  • We denote $\Pi=\Pi_{0}=\{x_{1}=0\}$ and, for $R>0$, $Z_{R}=\{x_{n+1}>R\}$.
  • For an arbitrary subset $A\subset\mathbb{R}^{n+1}$, we set
    $A_{+}(t)=\{x\in A:\mathbf{p}(x)\geq t\},\qquad A_{-}(t)=\{x\in A:\mathbf{p}(x)\leq t\},\qquad\delta_{t}(A)=A\cap\Pi_{t}.$
  • The right and left reflections of $A$ across $\Pi_{t}$ are
    $A_{+}^{*}(t)$ $=\{(2t-x_{1},x_{2},\ldots,x_{n+1}):(x_{1},\ldots,x_{n+1})\in A_{+}(t)\},$
    $A_{-}^{*}(t)$ $=\{(2t-x_{1},x_{2},\ldots,x_{n+1}):(x_{1},\ldots,x_{n+1})\in A_{-}(t)\}.$
  • We let $\pi:\mathbb{R}^{n+1}\to\Pi=\{x_{1}=0\}$ be the orthogonal projection
    $\pi(x_{1},\ldots,x_{n+1})=(0,x_{2},\ldots,x_{n+1}).$
Definition 6.3.
Let $A,B\subset\mathbb{R}^{n+1}$. We say that $A$ is on the right side of $B$ (and write $B\leq A$) if, for every $t\in\mathbb{R}$ and every $x\in\Pi_{t}$ such that
$\pi^{-1}(\{x\})\cap A\neq\emptyset\quad\text{and}\quad\pi^{-1}(\{x\})\cap B\neq\emptyset,$
we have
(20) $\sup\{\mathbf{p}(p):p\in\pi^{-1}(\{x\})\cap B\}\ \leq\ \inf\{\mathbf{p}(p):p\in\pi^{-1}(\{x\})\cap A\}.$
Remark 6.4.
When $A$ and $B$ are graphs over the same slice $\Pi_{t}$, the relation $B\leq A$ defines a partial order.
Now, in the following two lemmas we show that the moving-plane method applies to strictly convex entire translators that are smoothly asymptotic to the bowl-type solution.
Lemma 6.5.
Let $\gamma:\overline{\Gamma}\to[0,\infty)$ be a $1$-nondegenerate, $\alpha$-homogeneous, positive curvature function, and let $\Sigma=\operatorname{Graph}(u)$ be a strictly convex smooth $\gamma$-translator that is smoothly asymptotic to the bowl-type solution. Then, for every $R\gg 1$ there exists $r_{0}\geq R$ such that, for every $t>r_{0}$, the set $\Sigma_{+}(t)=\Sigma\cap\{x_{1}\geq t\}$ is the graph of a function over $\Pi_{t}$.
Proof.
First, from Section 4, the asymptotic behavior of the bowl-type solution $u_{B}:\mathbb{R}^{n}\to\mathbb{R}$ is given by
$u_{B}(x)=\begin{cases}\displaystyle\frac{|x|^{\alpha+1}}{\alpha+1}+\frac{a}{(\alpha-1)|x|^{\alpha-1}}+o\bigl(|x|^{1-3\alpha}\bigr),&\alpha\neq 1,\\[6.0pt] \displaystyle\frac{|x|^{2}}{2}-a\ln|x|+o\bigl(|x|^{-2}\bigr),&\alpha=1,\end{cases}$
where $a=\alpha\bigl(\partial_{x}\tilde{\gamma}(0,1)-\beta\bigr)\in\mathbb{R}$ [^6] When $\alpha=1$, $a=\partial_{x}\tilde{\gamma}(0,1)$.. Then, for any $C>0$ there exists $R\gg 1$ such that
$\partial_{x_{1}}u_{B}(x)\geq\begin{cases}\left(|x|^{\alpha-1}-\dfrac{|a|}{|x|^{\alpha+1}}-\dfrac{C}{|x|^{3\alpha+1}}\right)x_{1},&\alpha\neq 1,\\[6.0pt] \left(1-\dfrac{a}{|x|^{2}}-\dfrac{C}{|x|^{4}}\right)x_{1},&\alpha=1,\end{cases}$
whenever $|x|\geq R$. If $\alpha\geq 1$, we fix $\varepsilon_{0}>0$ and choose $r_{0}>R$ large enough such that, for all $r\geq r_{0}$,
$\begin{cases}r^{\alpha-1}-|a|r^{-\alpha-1}-Cr^{-(3\alpha+1)}\geq\dfrac{\varepsilon_{0}}{r_{0}},&\alpha>1,\\[6.0pt] 1-\dfrac{a}{r^{2}}-\dfrac{C}{r^{4}}\geq\dfrac{\varepsilon_{0}}{r_{0}},&\alpha=1.\end{cases}$
Then, for $x_{1}\geq r_{0}$ and $|x|\geq R$, we obtain $\partial_{x_{1}}u_{B}(x)\geq\varepsilon_{0}$.
On the other hand, if $\alpha\in\left(\frac{1}{3},1\right)$, we choose $r_{0}>R$ such that, for every $r\geq r_{0}$,
$r^{\alpha-1}-\dfrac{|a|}{r^{\alpha+1}}-\dfrac{C}{r^{1+3\alpha}}>0.$
It then follows that $\Sigma_{t}$ for $|x|\geq r_{0}$ can be expressed as the graph of a function $x_{1}=f(x_{2},\ldots,x_{n+1})$; otherwise, there exist $R\leq a<b$ such that
$0$ $=u(b,x_{2},\ldots,x_{n})-u(a,x_{2},\ldots,x_{n})$
$=\partial_{x_{1}}u(\xi,x_{2},\ldots,x_{n})(b-a)$
$\geq\left(r^{\alpha-1}(\xi)-\dfrac{|a|}{r^{\alpha+1}(\xi)}-\dfrac{C}{r^{1+3\alpha}(\xi)}\right)\frac{\xi(b-a)}{r(\xi)}$
$>0,$
where $r(\xi)=\sqrt{\xi^{2}+x_{2}^{2}+\ldots+x_{n}^{2}}>r_{0}$ and $\xi\in(a,b)$, a contradiction. Consequently, since the orthogonal projection along $\vec{e}_{1}$ preserves strict convexity of the sections, the global implicit function theorem in [global_implicit_function_theorem] implies that $\Sigma_{+}(t)$ is a graph over $\Pi_{t}$ for all $t>r_{0}$. ∎
Next, we show that, for $t$ large, the reflection of $\Sigma_{+}(t)$ across $\Pi_{t}$ lies to the right of the remaining part $\Sigma_{-}(t)$.
Lemma 6.6.
Under the hypotheses of Lemma 6.5, let $\varepsilon,\delta>0$. Then there exists $r_{1}=r_{1}(\varepsilon,\delta)>r_{0}$ such that
$u^{*}_{\Sigma,t}(x)-u_{\Sigma}(x)\geq\delta,$
for all $|x|\geq r_{1}$ and all $t>x_{1}+\varepsilon$, where
$u^{*}_{\Sigma,t}(x)=u_{\Sigma}(2t-x_{1},x_{2},\ldots,x_{n}).$
Proof.
Since $\Sigma$ is smoothly asymptotic to the bowl-type solution $u_{B}$, it is enough for us to carry out the estimates for $u_{B}$ and then absorb the $o(\cdot)$-terms into the choice of $r_{1}$. Indeed, we fix $\varepsilon,\delta>0$ and consider a point $x=(x_{1},x_{2},\ldots,x_{n})$ with $|x|\geq r$ for some $r\geq r_{0}$ (to be chosen), and a plane $\Pi_{t}$ with $t>x_{1}+\varepsilon$.
Then, we set $x^{\prime}=(2t-x_{1},x_{2},\ldots,x_{n})$, so that $x^{\prime}$ is the reflection of $x$ across $\Pi_{t}$. Notice that
$|x^{\prime}|^{2}=(2t-x_{1})^{2}+x_{2}^{2}+\cdots+x_{n}^{2}=4t(t-x_{1})+|x|^{2}.$
For convenience, we define $A=|x|^{2}$ and $B=4t(t-x_{1})\geq 4t\varepsilon>0$, so that $|x^{\prime}|^{2}=A+B$. Then, from (11), it follows that
$u_{B}(x)$ $=\frac{A^{\frac{\alpha+1}{2}}}{\alpha+1}+\frac{a}{(\alpha-1)A^{\frac{\alpha-1}{2}}}+o\bigl(A^{\frac{1-3\alpha}{2}}\bigr),$
$u_{B}(x^{\prime})$ $=\frac{(A+B)^{\frac{\alpha+1}{2}}}{\alpha+1}+\frac{a}{(\alpha-1)(A+B)^{\frac{\alpha-1}{2}}}+o\bigl((A+B)^{\frac{1-3\alpha}{2}}\bigr).$
Hence the difference satisfies
$u_{B}(x^{\prime})-u_{B}(x)$ $=\frac{(A+B)^{\frac{\alpha+1}{2}}-A^{\frac{\alpha+1}{2}}}{\alpha+1}+\frac{a}{\alpha-1}\left(\frac{1}{(A+B)^{\frac{\alpha-1}{2}}}-\frac{1}{A^{\frac{\alpha-1}{2}}}\right)+o\bigl(A^{\frac{1-3\alpha}{2}}\bigr).$
We now estimate $u_{B}(x^{\prime})-u_{B}(x)$ according to the values of $\alpha$:
  1. (1)
    Case $\alpha>1$. Since $A,B>0$ and $\frac{\alpha+1}{2}>1$, we have $(A+B)^{\frac{\alpha+1}{2}}-A^{\frac{\alpha+1}{2}}\geq B^{\frac{\alpha+1}{2}}$. Therefore
    $\frac{(A+B)^{\frac{\alpha+1}{2}}-A^{\frac{\alpha+1}{2}}}{\alpha+1}\ \geq\ \frac{B^{\frac{\alpha+1}{2}}}{\alpha+1}.$
    For the second term, we use the mean value theorem:
    $\frac{1}{(A+B)^{\frac{\alpha-1}{2}}}-\frac{1}{A^{\frac{\alpha-1}{2}}}=-\frac{\alpha-1}{2}\,\frac{B}{\xi^{\frac{\alpha+1}{2}}}$
    for some $\xi\in(A,A+B)$. Hence
    $\frac{a}{\alpha-1}\left(\frac{1}{(A+B)^{\frac{\alpha-1}{2}}}-\frac{1}{A^{\frac{\alpha-1}{2}}}\right)=-\frac{a}{2}\,\frac{B}{\xi^{\frac{\alpha+1}{2}}}.$
    Since $\xi\geq A=|x|^{2}$, we have
    $\left|\frac{a}{2}\,\frac{B}{\xi^{\frac{\alpha+1}{2}}}\right|\leq\frac{|a|}{2}\,\frac{B}{|x|^{\alpha+1}}.$
    Collecting terms, we obtain
    $u_{B}(x^{\prime})-u_{B}(x)\ \geq\ \frac{B^{\frac{\alpha+1}{2}}}{\alpha+1}-\frac{|a|}{2}\,\frac{B}{|x|^{\alpha+1}}-\frac{C}{|x|^{\alpha-1}},$
    for some $C>0$ coming from the $o(\cdot)$-term. For $t>x_{1}+\varepsilon$, we have
    $B=4t(t-x_{1})\geq 4t\varepsilon.$
    We first choose $t\geq r_{0}$ (from Lemma 6.5) and then choose $|x|\geq r_{1}$ so large that
    $\frac{(4t\varepsilon)^{\frac{\alpha+1}{2}}}{\alpha+1}-\frac{|a|}{2}\,\frac{4t\varepsilon}{r_{1}^{\alpha+1}}-\frac{C}{r_{1}^{\alpha-1}}\ \geq\ \delta.$
  2. (2)
    Case $\alpha=1$. Now
    $u_{B}(x)=\frac{A}{2}-a\ln(\sqrt{A})+o(A^{-1})=\frac{A}{2}-\frac{a}{2}\ln A+o(A^{-1}),$
    and similarly
    $u_{B}(x^{\prime})=\frac{A+B}{2}-\frac{a}{2}\ln(A+B)+o\bigl((A+B)^{-1}\bigr).$
    Thus
    $u_{B}(x^{\prime})-u_{B}(x)=\frac{B}{2}-\frac{a}{2}\ln\left(1+\frac{B}{A}\right)+o(A^{-1}).$
    Since $\ln(1+z)\leq z$ for $z>0$,
    $\frac{a}{2}\ln\left(1+\frac{B}{A}\right)\geq-\frac{|a|}{2}\,\frac{B}{A}.$
    Therefore,
    $u_{B}(x^{\prime})-u_{B}(x)\geq\frac{B}{2}-\frac{|a|}{2}\,\frac{B}{A}-\frac{C}{A}.$
    Recalling $B\geq 4t\varepsilon$, we fix $t\geq r_{0}$ and choose $|x|\geq r_{1}$ such that
    $2t\varepsilon-\frac{|a|}{2}\,\frac{4t\varepsilon}{r_{1}^{2}}-\frac{C}{r_{1}^{2}}\geq\delta.$
  3. (3)
    Case $\alpha\in\left(\frac{1}{3},1\right)$. We write
    $u_{B}(x^{\prime})-u_{B}(x)$ $\geq\dfrac{(A+B)^{\frac{\alpha+1}{2}}-A^{\frac{\alpha+1}{2}}}{\alpha+1}-\dfrac{a}{1-\alpha}\left((A+B)^{\frac{1-\alpha}{2}}-A^{\frac{1-\alpha}{2}}\right)$
    $=A^{\frac{1-\alpha}{2}}\left[\left(1+\frac{B}{A}\right)^{\frac{1-\alpha}{2}}\left(\frac{(A+B)^{\alpha}}{\alpha+1}-\frac{a}{1-\alpha}\right)-\left(A^{\alpha}-\frac{a}{1-\alpha}\right)\right].$
    We then choose $r_{1}=r_{1}(\delta,\varepsilon,r_{0},a,C)$ large enough such that, for any $r\geq r_{1}$,
    $\dfrac{r^{1-\alpha}}{\alpha+1}\bigl((r^{2}+4t\varepsilon)^{\alpha}-r^{2\alpha}\bigr)\geq\delta$
    holds.
Consequently, since $u_{\Sigma}-u_{B}=o(1)$ together with their first derivatives as $|x|\to\infty$, in all three cases we obtain
$u^{*}_{\Sigma,t}(x)-u_{\Sigma}(x)\geq\delta$
for all $|x|\geq r_{1}$ and $t>x_{1}+\varepsilon$.
From Lemma 6.6, letting $\delta\to 0$, it follows that for every $\varepsilon>0$ and every $t\geq r_{1}(\varepsilon)>r_{0}$ we have
$\Sigma_{-}(t+\varepsilon)\cap\{x_{1}\leq t\}\ \leq\ \Sigma_{+}^{*}(t+\varepsilon)\cap\{x_{1}\leq t\}.$
Moreover, since $\Sigma_{+}(t)$ is a graph over $\Pi_{t}$ for $t\geq r_{1}(\varepsilon)$, we also have
$\Sigma_{-}(t+\varepsilon)\cap\{t\leq x_{1}\leq t+\varepsilon\}\ \leq\ \Sigma_{+}^{*}(t+\varepsilon)\cap\{t\leq x_{1}\leq t+\varepsilon\}.$
We now define the moving-plane parameter set
$\mathcal{A}=\bigl\{\,t\in[0,\infty):\Sigma_{+}(t)\text{ is a graph over }\Pi\text{ and }\Sigma_{-}(t)\leq\Sigma_{+}^{*}(t)\,\bigr\}.$
By the previous discussion, $\mathcal{A}\neq\emptyset$, since $t+\varepsilon\in\mathcal{A}$ for all $t\geq r_{1}(\varepsilon)$.
Proof of Theorem 6.1.
By Lemmas 6.5 and 6.6, the set
$\mathcal{A}=\{t\in[0,\infty):\Sigma_{+}(t)\text{ is a graph over }\Pi\text{ and }\Sigma_{-}(t)\leq\Sigma_{+}^{*}(t)\}$
is nonempty. Using the standard Alexandrov moving-plane argument (compare with [Paco_2014, Thm. A]), we show that $\mathcal{A}$ is both open and closed in $0,\infty)$. Since $0,\infty)$ is connected, it follows that $\mathcal{A}=[0,\infty)$; in particular $0\in\mathcal{A}$.
Thus,
$\Sigma_{-}(0)\leq\Sigma_{+}^{*}(0)\quad\text{and, by symmetry,}\quad\Sigma_{-}^{*}(0)\leq\Sigma_{+}(0).$
Combining these two inequalities yields that $\Sigma$ is symmetric with respect to $\Pi=\{x_{1}=0\}$. By repeating the argument with planes orthogonal to other horizontal directions, and using that equation (2) is invariant under rotations fixing the $x_{n+1}$-axis, we conclude that $\Sigma$ is rotationally symmetric. This proves the theorem. ∎
Using similar comparison arguments, we obtain a growth estimate for entire convex solutions.
Corollary 6.7.
Let $\gamma:\overline{\Gamma}\to[0,\infty)$ be a positive, $1$-nondegenerate, normalized curvature function, and let $\Sigma=\{(x,u(x)):x\in\mathbb{R}^{n}\}$ be an entire, complete, convex $\gamma$-translator with polynomial growth
$C_{1}|x|^{a}\ \leq\ u(x)\ \leq\ C_{2}|x|^{b}\qquad\text{for }|x|\geq R,$
for some positive constants $a,b,C_{1},C_{2},R$. Then $a\leq\alpha+1\leq b$.
Proof.
Let $B$ denote the bowl-type translator in $\mathbb{R}^{n+1}$ corresponding to the same $1$-nondegenerate, positive, normalized curvature function $\gamma$. From the asymptotic expansion of $B$,
$u_{B}(x)=\begin{cases}|x|^{\alpha+1}+o\bigl(|x|^{1-3\alpha}\bigr),&\alpha\neq 1,\\ |x|^{2}-o(\ln|x|),&\alpha=1,\end{cases}$
we see that if $a>\alpha+1$, then for $|x|$ large enough the graph of $\Sigma$ lies strictly above a vertical translate of $B$. By strict convexity, we shift $\Sigma$ downward until first contact. However, the tangency principle (Theorem 6.2) implies that $\Sigma$ must coincide with a translate of $B$, contradicting $a>\alpha+1$. The case $b<\alpha+1$ is analogous, placing $B$ above $\Sigma$ and moving $\Sigma$ upward to first contact.
Corollary 6.8.
Let $\gamma:\overline{\Gamma}\to[0,\infty)$ be a signed $\alpha$-homogeneous curvature function that is $1$-nondegenerate and (n1)(n-1)-nondegenerate, with $\gamma(0,\ldots,0,1)=1$ and a continuous extension satisfying $\gamma(0)=0$. Then no strictly convex translator $\Sigma$ is contained in a round cylinder $\mathbb{S}^{n-1}(r)\times\mathbb{R}$ in $\mathbb{R}^{n+1}$.
Proof.
Given Theorem [santaella2025maximum, Thm. 1.3], which implies that no translator can be contained in a non-vertical round cylinder, it only remains for us to verify that no translator can be contained in a vertical one. We assume, by contradiction, that there exists a translator $\Sigma$ with these properties. Then, we choose $t_{0}\gg 1$ such that $\Sigma+t_{0}\vec{e}_{n+1}$ does not touch the catenoidal translator $W_{R}$, for $R\in(0,r)$.
We next translate $\Sigma+t_{0}\vec{e}_{n+1}$ downward, such that $W_{R}$ and $\Sigma-t_{0}\vec{e}_{n+1}$ have a first point of contact, see Fig. 4. This point is located either on the upper branch or on the lower branch of $W_{R}$. However, $\Sigma$ must then coincide with a translation of the bowl-type solution, which is impossible because it is contained in a round cylinder, finishing the proof. ∎
Figure 4. First contact point of $W_{R}$ and $\Sigma-t_{1}\vec{e}_{n+1}$. Figure prepared by Ignacio McManus.
7. Conclusions
In this work, we extend the theory of translating solitons in euclidean spaces for the mean curvature flow to a broad class of curvature functions that are symmetric, homogeneous, and monotone, under natural (non)degeneracy assumptions. The main novelty is the construction of catenoidal examples whose curvature functions fail to be continuous at the origin, showing that this phenomenon is compatible with the geometric structure of the corresponding solitons.
These constructions provide a framework for constructing $\Delta$-wing solitons for non-degenerate (n1)(n-1) normalized curvature functions, following the variational approaches employed by Hoffman-Ilmanen-Martin-White [hoffman2019graphical] as well as the more EDP-based techniques of Bourni-Langford-Tinaglia [bourni2020existence] under the condition of being rotationally symmetric at infinity, since such solitons arise naturally from the process of desingularizations between bowl-type translators and grim-reaper cylinders. Furthermore, the catenoidal translators we construct should allow stability results for complete graphical solutions of the associated $\gamma$-flow, in the spirit of Clutterbuck-Schnürer-Schulze [clutterbuck2007stability].